### v2

# Safe Output Regulation of Coupled Hyperbolic PDE-ODE Systems

Abstract
This paper presents a safe output regulation control strategy for a class of systems modeled by a coupled $2\times 2$ hyperbolic PDE-ODE structure, subject to fully distributed disturbances throughout the system. A state-feedback controller is developed by the nonovershooting backstepping method to simultaneously achieve exponential output regulation and enforce safety constraints on the regulated output that is the state furthest from the control input. To handle unmeasurable states and external disturbances, a state observer and a disturbance estimator are designed. Explicit bounds on the estimation errors are derived and used to construct a robust safe regulator that accounts for the uncertainties. The proposed control scheme guarantees that: 1) If the regulated output is initially within the safe region, it remains there; otherwise, it will be rescued to the safety within a prescribed time; 2) The output tracking error converges to zero exponentially; 3) The observer accurately estimates both the distributed states and external disturbances, with estimation errors converging to zero exponentially; 4) All signals in the closed-loop system remain bounded. The effectiveness of the proposed method is demonstrated through a UAV delivery scenario with a cable-suspended payload, where the payload is regulated to track a desired reference while avoiding collisions with barriers.
keywords:
hyperbolic PDEs; safe control; output regulation; boundary control
thanks: The material in this paper was not presented at any conference.
and
1 Introduction
1.1 Output Regulation of Coupled Hyperbolic PDEs
The output regulation problem seeks to design the control law that ensures the system output tracks desired references and/or rejects undesired disturbances. In recent years, considerable attention has been devoted to output regulation for infinite-dimensional systems governed by partial differential equations (PDEs). The Internal Model Principle (IMP), serving as a fundamental theoretical framework for output regulation, whose primary advantage lies in its inherent robustness against internal uncertainties, has been extensively investigated for infinite-dimensional systems in [18, 19, 20, 21].
In addition, the integration of output regulation theory with the backstepping methodology [14] has led to systematic and constructive procedures for regulator design in PDE systems. Upon the backstepping-based boundary stabilization of $2\times 2$ hyperbolic PDEs presented in [24] and [8], initial contributions toward backstepping-based output regulation for $2\times 2$ hyperbolic PDEs appeared in [1]. Subsequent extensions to more general heterodirectional transport PDEs were proposed in [2], [11], and [33]. For wave-type PDEs, similar regulator designs were developed in [34] and [35]. Research efforts were also directed at PDE-ODE interconnected systems, which naturally arise in many applications. Backstepping boundary control designs for interconnected $2\times 2$ hyperbolic PDEs and ODEs were proposed in [17], and the ODE-PDE-ODE configurations were further addressed in [28], [26], [9], and [7]. Subsequently, the output regulation of hyperbolic PDEs sandwiched between two ODEs was explored in [37], and the regulation problems involving coupled wave PDE-ODE systems and hyperbolic PDEs coupled with nonlinear ODEs were investigated in [12] and [36], respectively. A more detailed literature review of backstepping output regulation/disturbance rejection in PDEs can be found in Sec. 3.5 of [25]. Despite progress in PDE output regulation, guaranteed safety remains unaddressed. Ensuring that outputs remain within safe operating regions is crucial for autonomous systems, thus the problem of safe output regulation of PDEs is essential and open.
1.2 Safe Control
Control Barrier Functions (CBFs) have emerged as a powerful tool for safe control design, as introduced in [5], where system states are constrained to remain within a designated safe region by ensuring the non-negativity of an appropriately constructed CBF. To enforce this constraint within the closed-loop system, CBF-based control strategies are typically combined with a quadratic program (QP) safety filter, which overrides potentially unsafe nominal control inputs to generate safe control actions. High relative degree CBFs have been further developed in [22], [31], and [32], after the presence of non-overshooting control design techniques in [15]. Based on the tools developed in [15], mean-square stabilization of stochastic nonlinear systems toward an equilibrium on the barrier was achieved in [16], and prescribed-time safety (PTSf)—which guarantees safety over a user-specified finite time horizon—was proposed in [3].
However, safe control of PDE systems remains largely underexplored. The first result in this area was presented in [13], which proposed a CBF-based boundary control scheme for a Stefan model involving parabolic PDEs with actuator dynamics. Adaptive safe control for hyperbolic PDEs was proposed in [30]. Yet, these approaches address full-state control but do not address output regulation in the presence of unmeasured states and external disturbances.
1.3 Main Contributions
1) In contrast to traditional output regulation of hyperbolic PDEs [1], [11], [12], [37], our approach not only ensures that the system output tracks the desired trajectory but also guarantees that it stay within a prescribed safe region.
2) Compared to our previous work [30] on safe control of coupled hyperbolic PDEs, this paper presents several improvements: a) Beyond the state-feedback framework, a safe output-feedback controller is developed, which accounts for unmeasured PDE states and external disturbances during the safe regulation; b) Instead of restricting safety to positivity, a general barrier function $h$ is introduced, enabling broader and more flexible safety specifications; c) If the system starts outside the safe set, the proposed controller guarantees that it will be rescued to the safety within a prescribed time that can be selected as a design parameter by the user.
3) To the best of our knowledge, this work presents the first result on safe output regulation for PDE systems. We apply the control input to a UAV that transports a cable-suspended payload, successfully regulating the payload at the bottom of the cable to track a desired reference while avoiding collisions with surrounding obstacles.
1.4 Notation
  • The symbol $\mathbb{N}$ denotes the set $\{1,2,\cdots\}$.
  • We use the notation $L^{\infty}(0,1)$ for the standard Banach space of essentially bounded, measurable functions $f:(0,1)\to\mathbb{R}$, equipped with the norm $\|f\|_{\infty}:=\text{ess sup}_{x\in(0,1)}|f(x)|<+\infty$ for $f\in L^{\infty}(0,1)$.
  • Let $u:\mathbb{R}_{+}\times[0,1]\rightarrow\mathbb{R}$ be given. We use the notation $u[t]$ to denote the profile of $u$ at certain $t\geq 0$, i.e., $u[t]=u(x,t)$ for all $x\in[0,1]$.
  • Let $|Y|$ be Euclidean norm of the vector $Y$, and $\|A\|_{2}$ denotes the induced 2-norm of the matrix $A$.
  • We define $\underline{x}_{j}:=[x_{1},x_{2},\cdots,x_{j}]^{T}$.
For ease of presentation, we omit or simplify the arguments of functions and functionals when no confusion arises. Besides, if $a>b$ happens in $\sum_{i=a}^{b}$ of this paper, it means that the result is zero.
2 Problem Formulation
2.1 Plant
The considered plant is
$\dot{Y}(t)$ $=AY(t)+Bw(0,t)+G_{1}d(t),$ (1)
${z_{t}}(x,t)$ $=-q_{1}{z_{x}}(x,t)+{d_{1}}w(x,t)+G_{2}(x)d(t),$ (2)
${w_{t}}(x,t)$ $=q_{2}{w_{x}}(x,t)+{d_{2}}z(x,t)+G_{3}(x)d(t),$ (3)
$z(0,t)$ $=pw(0,t)+CY(t)+G_{4}d(t),$ (4)
$w(1,t)$ $=qz(1,t)+G_{5}d(t)+U(t),$ (5)
$\forall(x,t)\in[0,1]\times[0,\infty)$, where the function $U(t)$ is the control input to be designed. The scalars $z(x,t)\in\mathbb{R},w(x,t)\in\mathbb{R}$ are states of the PDEs, $Y^{T}(t)=[y_{1},y_{2},\cdots,y_{n}]\in\mathbb{R}^{n}$ are ODE states. The disturbances $d(t)\in\mathbb{R}^{m_{d}\times 1}$ are unmeasurable. The plant parameters $d_{1},d_{2}$, $q$, the transport speed $q_{1}>0,q_{2}>0$, the matrices $G_{1}\in\mathbb{R}^{n\times m_{d}},G_{2},G_{3}\in(C[0,1])^{1\times m_{d}}$ regarding the disturbance input locations, and $p\neq 0$, are known and arbitrary. The subsystem $Y$-ODE is strict-feedback linear, where the matrix $A$, the column vector $B$ are in the form of
$A=$ $\left(\begin{array}[]{cccccc}a_{1,1}&1&0&0&\cdots&0\\ a_{2,1}&a_{2,2}&1&0&\cdots&0\\ &&\vdots&&&\\ a_{n-1,1}&a_{n-1,2}&a_{n-1,3}&a_{n-1,4}&\cdots&1\\ a_{n,1}&a_{n,2}&a_{n,3}&a_{n,4}&\cdots&a_{n,n}\\ \end{array}\right),$ (11)
and $B=(0,0,\cdots,b)^{T}$, with arbitrary constants $a_{i,j}$, and $b>0$ (without any loss of generality for $b<0$). This indicates that the $Y$-ODE is in the controllable form, which covers many practical models. The matrix $C\in\mathbb{R}^{1\times n}$ is arbitrary.
2.2 Exogenous signal model
Like many articles on output regulation [11], the reference and disturbance signals are considered to be generated by a finite-dimensional exogenous model:
$\dot{v}(t)=Sv(t),~~t>t_{0},~~v(t_{0})\in\mathbb{R}^{n_{v}}$ (12)
where the spectrum $\sigma(S)$ of the known diagonalizable matrix $S\in\mathbb{R}^{n_{v}\times n_{v}}$ only contains eigenvalues on the imaginary axis, which allows the modelling of bounded and persistently acting exogenous signals. In particular, this signal model can generate the exogenous signals as constant or trigonometric functions of time as well as linear combinations of both signal forms. Defining $S={\rm bdiag}(S_{r},S_{d})$ with $S_{r}\in\mathbb{R}^{n_{r}\times n_{r}}$, $S_{d}\in\mathbb{R}^{n_{d}\times n_{d}}$, and $v={\rm col}(v_{r},v_{d})$ where $v_{r}\in\mathbb{R}^{n_{r}}$, $v_{d}\in\mathbb{R}^{n_{d}}$, one obtains the reference model and the disturbance model, respectively,
$\dot{v}_{r}(t)=S_{r}v_{r}(t),~\dot{v}_{d}(t)=S_{d}v_{d}(t)$ (13)
where $n_{r}+n_{d}=n_{v}$. Also, we have
$d(t)=P_{d}v(t)=\bar{P}_{d}v_{d}(t),~~P_{d}=\bar{P}_{d}E_{d},$ (14)
$r(t)=P_{r}v(t)=\bar{P}_{r}v_{r}(t),~~P_{r}=\bar{P}_{r}E_{r},$ (15)
where $P_{d}\in\mathbb{R}^{m_{d}\times n_{v}}$, $P_{r}\in\mathbb{R}^{1\times n_{v}}$, $\bar{P}_{d}\in\mathbb{R}^{m_{d}\times n_{d}}$, $\bar{P}_{r}\in\mathbb{R}^{1\times n_{r}}$, $E_{r}=[I_{r},0]\in\mathbb{R}^{n_{r}\times n_{v}}$, $E_{d}=[0,I_{d}]\in\mathbb{R}^{n_{d}\times n_{v}}$ where $I_{d}$, $I_{r}$ are identity matrices with dimensions of $n_{d}$, $n_{r}$, respectively.
Assumption 1.
The pairs $(S_{d},\bar{P}_{d})$ and $(S_{r},\bar{P}_{r})$ are observable.
2.3 Control Objective
There are two measurable outputs in the considered plant, one is $z(1,t)$ that will be utilized in the extended observer for both unmeasurable states and exogenous signals, and the other one is $y_{1}=C_{1}Y(t)$ that is the output to be safely regulated, which is described below. The vector $C_{1}=[1,0,\cdots,0]$ satisfies the following assumption.
Assumption 2.
The pair $(A,C_{1})$ is observable.
The control objective is to ensure that the output tracking error $e(t)=y_{1}(t)-r(t)$ of the system converges to zero while staying in the time-varying safe region in the sense that $e(t)\in\mathcal{C}(t),\forall t\geq\bar{t}_{0}$, where $\mathcal{C}(t)$ is a time-dependent safe set $\mathcal{C}(t):=\{e\in\mathbb{R}\mid h(e,t)\geq 0\}$ defined by the time-varying barrier function $h(e(t),t)$. Focusing on $y_{1}$ is only for ease of exposition and does not restrict the applicability of the result. The result is not applicable only to systems whose first state only is constrained. The result is applicable to all systems diffeomorphically transformable into the form (1), (11), which is equivalent to the input-output feedback linearization-like treatment with high-order CBFs (HOCBFs). And the result is also applicable, with an observer, to all systems transformable into (1) with a transformation acting not only on the measured state but on the entire (unmeasured) state. An example is given in the following remark.
Remark 1.
The control design is applicable to the case where the distal ODE subsystems (denoted as the XX–ODE) are described by general matrices $\bar{A}\in\mathbb{R}^{n},\bar{B}\in\mathbb{R}^{n\times 1},\bar{C}\in\mathbb{R}^{1\times n}$, assuming that the pair $(\bar{A},\bar{C})$ is observable and that the XX–ODE has relative degree $n$. Such systems are also covered by the framework considered in this paper. Indeed, the XX–ODE can be transformed into the YY–ODE studied here via the coordinate transformation $Y(t)=RX(t)$, $R=[\bar{C}^{T},(\bar{C}\bar{A})^{T},\ldots,(\bar{C}\bar{A}^{n-1})^{T}]^{T}$. Since $(\bar{A},\bar{C})$ is observable, the matrix $R$ is nonsingular. The relative degree $n$ that implies that $\bar{C}\bar{A}^{i}\bar{B}=0$ for $i=0,\cdots,n-2$ and $\bar{C}\bar{A}^{n-1}\bar{B}\neq 0$. Then it is ensured that the transformed system’s system matrix, which is in the companion (observable canonical) form whose last row consists of the coefficients of the characteristic polynomial of $\bar{A}$ according to Hamilton-Cayley theorem, is covered by the structure $A$ in (11), and the resulting input and output matrices also match $B,C_{1}$ matrices considered in this paper. Furthermore, the system output is preserved under the transformation, namely, $\bar{C}X=C_{1}Y=y_{1}$. Therefore, the safe regulation of the YY–ODE output achieved by the following control design directly implies the safe regulation of the original XX–ODE output.
The safety goal is now to ensure $h(e(t),t)\geq 0$ from the initial regulation time:
$\bar{t}_{0}=t_{0}+\frac{1}{q_{2}}.$ (16)
The introduction of initial regulation time $\bar{t}_{0}$ in (16) is due to the delay property of the transport PDE. Specifically, there is no control influence on the distal ODE over the time interval $t_{0},\bar{t}_{0}]$, where $t_{0}$ is the initial time. We impose the following assumption regarding the time-varying barrier function $h$ for the output signal.
Assumption 3.
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\{\ell\in\mathbb{R}|h(\ell,t)\geq 0\}$, $t\in[\bar{t}_{0},\infty)$ when $h$ is positive at $t=\bar{t}_{0}$, or otherwise $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[\bar{t}_{0},\infty)$. Besides, there exists a positive constant $\xi$ such that $|e(t)|<\xi|h(e(t),t)|$
Assumption 3 implies that $\lim_{t\to\infty}h(e(t),t)=0\Rightarrow\lim_{t\to\infty}e(t)=0$, which means the prescribed safe region converges to the desired reference as $t\to\infty$, which is reasonable in the output regulation problem.
In the derivations below, we work under the standing strong assumption that the plant, and later the corresponding closed-loop and observer systems, admit classical solutions on the time interval under consideration, so that all temporal and spatial derivatives appearing in the PDEs and all boundary traces used in the control law and proofs are well defined pointwise.
3 Nominal Safe Control Design
Applying (14), then (1)–(5) are rewritten as
$\dot{Y}(t)$ $=AY(t)+Bw(0,t)+\acute{G}_{1}v(t),$ (17)
${z_{t}}(x,t)$ $=-q_{1}{z_{x}}(x,t)+{d_{1}}w(x,t)+\acute{G}_{2}(x)v(t),$ (18)
${w_{t}}(x,t)$ $=q_{2}{w_{x}}(x,t)+{d_{2}}z(x,t)+\acute{G}_{3}(x)v(t),$ (19)
$z(0,t)$ $=pw(0,t)+CY(t)+\acute{G}_{4}v(t),$ (20)
$w(1,t)$ $=qz(1,t)+\acute{G}_{5}v(t)+U(t),$ (21)
where $\acute{G}_{i}=G_{i}P_{d},~i=1,\cdots,5$. The matrixes $\acute{G}_{i}\in\mathbb{R}^{1\times n_{v}},i=2,\cdots,5$ and $\acute{G}_{1}:=[\acute{g}_{1};\cdots;\acute{g}_{n}]\in\mathbb{R}^{n\times n_{v}}$ with $\acute{g}_{j}\in\mathbb{R}^{1\times n_{v}},j=1,\cdots,n$, are known.
3.1 First transformation
First, we propose the following transformation to derive and rewrite the tracking error dynamics in a form that will be helpful for the next transformation to barrier functions,
$Z(t)=T_{z}Y(t)+T_{v}v(t),$ (22)
where $Z(t)=[z_{1},\cdots,z_{n}]^{T}$ and the matrix $T_{z}\in\mathbb{R}^{n\times n}$ is
$T_{z}=\left(\begin{array}[]{ccccc}1&0&0&0&0\\ \varrho_{1,1}&1&0&0&0\\ \varrho_{2,1}&\varrho_{2,2}&1&0&0\\ \vdots\\ \varrho_{n-1,1}&\varrho_{n-1,2}&\cdots&\varrho_{n-1,n-1}&1\\ \end{array}\right)$ (28)
with the constants $\varrho_{i,j}$ defined by
$\varrho_{1,1}=a_{1,1},$ (29)
$\varrho_{2,1}=a_{2,1}+\varrho_{1,1}a_{1,1},~~\varrho_{2,2}=a_{2,2}+\varrho_{1,1}$ (30)
and, for $i=3,\cdots,n$, by
$\varrho_{i,1}=a_{i,1}+\sum_{j=1}^{i-1}\varrho_{i-1,j}a_{j,1},$ (31)
$\varrho_{i,\imath}=a_{i,\imath}+\varrho_{i-1,\imath-1}+\sum_{j=\imath}^{i-1}\varrho_{i-1,j}a_{j,\imath},~\forall\imath=2,\cdots,i-1,$ (32)
$\varrho_{i,i}=a_{i,i}+\varrho_{i-1,i-1},$ (33)
and where the matrix $T_{v}\in\mathbb{R}^{n\times n_{v}}$ is
$T_{v}=-[P_{r};\lambda_{1}+P_{r}S^{1};\cdots;\lambda_{n-1}+P_{r}S^{n-1}]$ (34)
with the vectors $\lambda_{i}$ defined by
$\lambda_{0}=0,$ (35)
$\lambda_{i}=-\sum_{j=1}^{i-1}\varrho_{i-1,j}\acute{g}_{j}-\acute{g}_{i}+\lambda_{i-1}S~~\forall~i=1,\cdots,n.$ (36)
Applying the transformation (22), we now convert the $Y$-ODE (17) into
$\dot{Z}(t)=A_{\rm z}Z(t)+Bw(0,t)+BK^{T}Y(t)-\bar{G}_{0}v(t)$ (37)
where $T_{z}B=B$ is recalled, and
$A_{\rm z}=\left(\begin{array}[]{ccccccc}0&1&0&0&\cdots&0&0\\ 0&0&1&0&\cdots&0&0\\ 0&0&0&1&\cdots&0&0\\ &&\vdots&&&&\\ 0&0&0&0&\cdots&0&1\\ 0&0&0&\cdots&0&0&0\\ \end{array}\right)$ (44)
and where
$K^{T}=\frac{1}{b}[\varrho_{n,1},\cdots,\varrho_{n,n}]_{1\times n},$ (45)
with $\bar{G}_{0}=\frac{1}{b}B(\lambda_{n}+P_{r}S^{n})$. Regarding $Z$-ODE in (37), $z_{1}$ is the tracking error that is to be regulated to zero, i.e.,
$z_{1}(t)=e(t)$ (46)
and the system matrix $A_{\rm z}$ represents a chain-of-integrators which will be helpful for advancing to high-relative-order barrier functions in the next transformation. The details of the first transformation are shown in Appendix A.
3.2 Second transformation
For the purpose of establishing barrier functions for the distal ODE, inspired by [3], we propose the following transformation
$h_{1}(z_{1}(t),t)=h(e(t),t)+\sigma(t),$ (47)
$h_{i}(\underline{z}_{i},t)=\sum_{\jmath=1}^{i-1}\frac{\partial h_{i-1}}{\partial z_{\jmath}}z_{\jmath+1}+\frac{\partial h_{i-1}}{\partial t}+k_{i-1}h_{i-1},$ (48)
for $i=2,3,\cdots,n$, with
$\sigma(t)=\begin{cases}\begin{cases}e^{\frac{1}{t_{a}^{2}}}({-h(e(\bar{t}_{0}),\bar{t}_{0})+\epsilon})e^{\frac{-1}{(t-\bar{t}_{0}-t_{a})^{2}}},t\in[t_{0},\bar{t}_{0}+t_{a}),\\ 0,~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~t\geq\bar{t}_{0}+t_{a}\end{cases}\\ if~~~h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0.\\ 0,~~t\geq t_{0},~~~~~if~~h(e(\bar{t}_{0}),\bar{t}_{0})>0\end{cases}$ (49)
where $t_{a}$ and $\epsilon$ are arbitrarily positive design parameters. The function $\sigma(t)$ is designed to address scenarios where the states are in the unsafe region at the initial regulation time $\bar{t}_{0}$. Specifically, when the initial state $e(\bar{t}_{0})$ fall outside the original safe region—indicated by the condition $(h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$—a new barrier function is created to guide the state back to the original safe region. The constant $\epsilon$ indicates the distance from the safe barrier of the initial value of the new barrier function $h_{1}$, and $t_{a}$ is the upper bound of the time taken to rescue to the safe region. We only ensure the state, which falls into the unsafe region at the initial regulation time $\bar{t}_{0}$, to rescue to safety within a prescribed time $t_{a}$, but not prescribed-time regulation to the safe boundary. Please note that even though $t_{a}$ can be assigned arbitrarily, a smaller value of $t_{a}$ leads to larger control gains, and thus a proper trade-off should be considered in practical implementation. The evolution of $e(t)$ on $t\in[t_{0},\bar{t}_{0}]$ can be expressed as the initial states as
$e(t_{0}+a)=C_{1}Z(t_{0}+a)$ (50)
$\forall a\in[0,\frac{1}{q_{2}}]$, where $Z(t_{0}+a)$ is defined by (B.5) in Appendix B, which is determined by the initial states $Y(t_{0}),w[t_{0}],z[t_{0}]$ and $v(t_{0})$. Therefore, $h(e(\bar{t}_{0}),\bar{t}_{0})$ can be determined by the initial states of the plant according to (50), denoted as
$h(e(\bar{t}_{0}),\bar{t}_{0})=\mathcal{P}(Y(t_{0}),z[t_{0}],w[t_{0}],v(t_{0}))$ (51)
where the function $\mathcal{P}$ is determined by inserting (50) with $a=\frac{1}{q_{2}}$, i.e., $e(\bar{t}_{0})$, into the barrier function $h(e,t)$. Note that $\sigma(t)$ is continuous and has continuous derivatives of all orders. Therefore, together with Assumption 3, $h_{1}(z_{1}(t),t)$ is $n$ times differentiable with respect to each of its arguments, i.e., $z_{1}$ and $t$. Please also note that $\dot{h}(e(t),t)$ denotes full derivative with respect to $t$ whose calculation using the chain rule, and $\frac{\partial h(e(t),t)}{\partial t}$ denotes partial derivative with respect to one of its variables $t$. Define
$\frac{\partial h(e(t),t)}{\partial e(t)}:=\vartheta(z_{1}(t),t),$ (52)
because of $z_{1}(t)=e(t)$ (46). Considering the fact
$\frac{\partial h_{n}}{\partial z_{n}}=\frac{\partial h_{n-1}}{\partial z_{n-1}}=\frac{\partial h_{n-2}}{\partial z_{n-2}}=\cdots=\frac{\partial h_{1}}{\partial z_{1}}=\frac{\partial h}{\partial e}=\vartheta$ (53)
that is obtained from (48), using (48) for $i=n$, we have $\dot{h}_{n}(\underline{z}_{n},t)+k_{n}h_{n}=b(\vartheta w(0,t)+f(\underline{y}_{n}(t),v(t),t)-\vartheta\frac{1}{b}(\lambda_{n}+P_{r}S^{n})v(t)+\vartheta K^{T}Y(t))$, where
$f(\underline{z}_{n}(t),t)=\frac{1}{b}\bigg(\sum_{\jmath=1}^{n-1}\frac{\partial h_{n}}{\partial z_{\jmath}}z_{\jmath+1}+\frac{\partial h_{n}}{\partial t}$
$+k_{n}[\sum_{\jmath=1}^{n-1}\frac{\partial h_{n-1}}{\partial z_{\jmath}}z_{\jmath+1}+\frac{\partial h_{n-1}}{\partial t}+k_{n-1}h_{n-1}]\bigg).$ (54)
Applying (47), (48) to convert (37) into
$\dot{H}(t)=$ $A_{\rm h}H(t)+B\bigg(f(\underline{z}_{n}(t),t)+\vartheta w(0,t)$
$-\vartheta\frac{1}{b}(\lambda_{n}+P_{r}S^{n})v(t)+\vartheta K^{T}Y(t)\bigg)$ (55)
where $H=[h_{1},\cdots,h_{n}]^{T}$, and where
$A_{\rm h}=\left(\begin{array}[]{ccccccc}-k_{1}&1&0&0&\cdots&0&0\\ 0&-k_{2}&1&0&\cdots&0&0\\ 0&0&-k_{3}&1&\cdots&0&0\\ &&\vdots&&&&\\ 0&0&0&0&\cdots&0&1\\ 0&0&0&\cdots&0&0&-k_{n}\\ \end{array}\right).$ (62)
This chain structure $\dot{H}(t)=A_{\rm h}H(t)$, also known as the form of high relative-degree CBFs [22], [31], was proposed in [15] for the nonovershooting control.
3.3 Third transformation
In order to remove the in-domain coupling destabilizing terms from the $2\times 2$ hyperbolic PDE system, the extra terms in (37), and the dependence on the signal model state $v(t)$, we propose the following backstepping transformations:
$\alpha(x,t)=z(x,t)-\int_{0}^{x}{\phi}(x,y)z(y,t)dy$
$-\int_{0}^{x}{\varphi}(x,y)w(y,t)dy-\gamma(x)Y(t)-\bar{\gamma}(x)v(t)-\bar{\varsigma}(x,t),$ (63)
$\beta(x,t)=w(x,t)-\int_{0}^{x}{\Psi}(x,y)z(y,t)dy$
$-\int_{0}^{x}{\Phi}(x,y)w(y,t)dy-\lambda(x)Y(t)-\bar{\lambda}(x)v(t)-{\varsigma}(x,t)$ (64)
where ${\phi},\varphi,\gamma,\bar{\gamma},\Psi,\Phi,\lambda,\bar{\lambda}$ are defined in Appendix C. In particular, equations (C.12)–(C.15) for $\bar{\gamma},\bar{\lambda}$ are the solvable regulator equations because the output $y_{1}$ is a flat output of the considered plant (1)–(5). Besides, the functions ${\varsigma}(x,t)$ and $\bar{\varsigma}(x,t)$ satisfy
$\varsigma_{t}(x,t)=q_{2}{\varsigma}_{x}(x,t),~~\varsigma(0,t)=\frac{-1}{\vartheta(z_{1}(t),t)}f(\underline{z}_{n},t),$ (65)
$\bar{\varsigma}_{t}(x,t)=-q_{1}\bar{\varsigma}_{x}(x,t),~~\bar{\varsigma}(0,t)=p*\varsigma(0,t).$ (66)
The solution $\varsigma(x,t)$ is
$\varsigma(x,t)=\frac{-1}{\vartheta(z_{1}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}})}f\left(\underline{z}_{n}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}}\right)$
where $\vartheta(z_{1}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}})$ is nonzero according to Assumption 3. The solution (3.3) is determined by the prediction $Z(t+\frac{x}{q_{2}})$ given by (B.5) that is expressed by the current states $Y(t)$, $w[t]$, $z[t]$ and $v(t)$ recalling (22) to replace $Z(t)$ with $Y(t)$. By (63), (64), the system (2)–(5) with (37) is converted into
$\dot{H}(t)=$ $A_{\rm h}H(t)+B\vartheta(z_{1}(t),t)\beta(0,t),$ (68)
${\alpha_{t}}(x,t)=$ $-{q_{1}}{\alpha_{x}}(x,t),~~{\beta_{t}}(x,t)={q_{2}}{\beta_{x}}(x,t),$ (69)
$\alpha(0,t)=$ $p\beta(0,t),~~\beta(1,t)=0$ (70)
by choosing the control input as
$U(t)=$ $-qz(1,t)+\int_{0}^{1}{\Psi}(1,y)z(y,t)dy$
$+\int_{0}^{1}{\Phi}(1,y)w(y,t)dy+\lambda(1)Y(t)$
$-(\acute{G}_{5}-\bar{\lambda}(1)){v}(t)+\varsigma(1,t)$ (71)
where $\varsigma(1,t)$ is given by (47)–(49), (54), (3.3), (B.5), ${\Psi}(1,y),{\Phi}(1,y)$, $\lambda(1)$ are determined by (C.16)–(C.21), and $\bar{\lambda}(1)$ is given by (C.12), (C.13), that is,
$\varsigma(1,t)=$ $\frac{-f(\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}{\vartheta(C_{1}\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}$
$:=$ $g(Z(t),z[t],w[t],v(t),t)$
$=$ $g(T_{z}Y(t)+T_{v}v(t),z[t],w[t],v(t),t).$ (72)
Remark 2.
If actuator dynamics are included, the proposed safe output regulation scheme can be extended to an ODE–PDE–ODE configuration using the approach in [30]. It can also accommodate uncertain input delays in hyperbolic PDE–ODE cascades by integrating the method in [29].
3.4 Selection of safe design parameters
The design parameters $k_{i}$, $i=1,2,\cdots,n-1$ satisfy
$k_{i}>\max\{0,\acute{k}_{i}\}$ (73)
where $\acute{k}_{i}(v(t_{0}))=\frac{-1}{h_{i}(\underline{z}_{i}(\bar{t}_{0}),\bar{t}_{0})}\big(\sum_{\jmath=1}^{i}\frac{\partial h_{i}}{\partial z_{\jmath}}z_{\jmath+1}(\bar{t}_{0})+\frac{\partial h_{i}}{\partial t}\big)$. By (B.5) and (22), $Z(\bar{t}_{0})$ can be expressed as the initial states $z[t_{0}],w[t_{0}],Y(t_{0})$ and $v(t_{0})$. We write $\acute{k}_{i}(v(t_{0}))$ to emphasize that it depends on the unknown initial values $v(t_{0})$ of the external signal, which will be dealt with in the next section. The gain condition (73) for the nonovershooting control ensures the following lemma.
Lemma 1.
With the design parameters $\kappa_{i}$, $i=1,\cdots,n-1$ satisfying (73), the high-relative-degree ODE CBFs is initialized positively, i.e., $h_{i}(\underline{z}_{i}(\bar{t}_{0}),\bar{t}_{0})>0$, $i=1,\cdots,n$.
Proof.
According to (47) and (49), we know $h_{1}(z_{1}(\bar{t}_{0}),\bar{t}_{0})>0$. Recalling (48), we have $h_{i}(\underline{z}_{i}(\bar{t}_{0}),\bar{t}_{0})=\sum_{\jmath=1}^{i-1}\frac{\partial h_{i-1}}{\partial z_{\jmath}}z_{\jmath+1}(\bar{t}_{0})+\frac{\partial h_{i-1}}{\partial t}+k_{i-1}h_{i-1}(\underline{z}_{i-1}(\bar{t}_{0}),\bar{t}_{0})$, $i=2,3,\cdots,n$. The lemma is thus obtained recalling (73).
We give the following definition about the safe initial condition, which is the sufficient and necessary condition under which the output state stays in the safe region before the control input arrives at the distal ODE subsystem. It is the counterpart to the initial state restriction, i.e., Assumption 1, in the work [4] about safe control with delays.
Definition 1.
The safe initial condition is defined as $\wp(Y(t_{0}),z[t_{0}],w[t_{0}],v(t_{0}),a)\geq 0$ for $a\in[0,\frac{1}{q_{2}}]$ and $\wp(Y(t_{0}),z[t_{0}],w[t_{0}],v(t_{0}),\frac{1}{q_{2}})\neq 0$, where the function $\wp$ is given by (B.6).
This definition is the sufficient and necessary condition of $h(e(t),t)\geq 0$ for $t\in[t_{0},\bar{t}_{0}]$, i.e., the safety holds for the uncontrolled period at the beginning, and $h(e(\bar{t}_{0}),\bar{t}_{0})\neq 0$ which means that the state does not stay at the safe boundary at the initial regulation time $t=\bar{t}_{0}$ for the distal ODE.
3.5 Result of the state-feedback safe control
Theorem 1.
Assume that the closed-loop system consisting of (1)–(5) and (71) admits a classical solution. For initial data $w[t_{0}]\in L^{\infty}(0,1)$, $z[t_{0}]\in L^{\infty}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, and for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$, satisfying (73), the following properties hold:
1) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
2) Safety is ensured in the sense that:
a) When the safe initial condition in Definition 1 is satisfied, the safety is ensured from $t=t_{0}$, i.e., $h(e(t),t)\geq 0,~\forall t\geq t_{0}$;
b) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})>0$, the safety is ensured from $t=\bar{t}_{0}$, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}$;
c) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, the state will return and stay in the safe region no latter than a finite time $\bar{t}_{0}+t_{a}$ where a constant $t_{a}>0$ can be arbitrarily assigned by users, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}+t_{a}$.
Proof.
1) Recalling the target system (68)–(70) and Assumption 3 indicates that $\vartheta$ is bounded, by the method of characteristics, it is obtained that $\beta[t],\alpha[t],H(t)$ are bounded all the time, and moreover, $\beta[t]\equiv 0$, $\alpha[t]\equiv 0$ for $t\geq\frac{1}{q_{1}}+\frac{1}{q_{2}}$ and $|H(t)|$ is exponentially convergent to zero recalling the fact that $A_{\rm h}$ defined in (62) is Hurwitz.
We then obtain from (47), (49), and Assumption 3 that $z_{1}$ is convergent to zero. According to (48), we have $\frac{\partial h_{i}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+\frac{\partial h_{i-1}^{2}}{\partial z_{i-1}\partial t}+k_{i-1}\frac{\partial h_{i-1}}{\partial z_{i-1}}=\frac{\partial h_{i-1}}{\partial z_{i-2}}+k_{i-1}\frac{\partial h_{1}}{\partial z_{1}}+\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}=\frac{\partial h_{2}}{\partial z_{1}}+\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=2}^{j=i-1}k_{j}+(i-2)\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}=\frac{\partial h_{1}}{\partial z_{1}}\sum_{j=1}^{j=i-1}k_{j}+(i-1)\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t},~~\forall i\geq 2,$ where (53) has been recalled. Repeating the same process for $\frac{\partial h_{i}}{\partial z_{i-j}}$, $j+1\leq i\leq n$, $j=2,\cdots,n-1$, we obtain $\sum_{i=1}^{n}\sum_{j=1}^{i}|\frac{\partial h_{i}}{\partial z_{j}}|\leq\mathcal{M}\big(|\frac{\partial h_{1}}{\partial z_{1}}|+\sum_{j=1}^{n-1}|\frac{\partial h_{1}^{n}}{\partial z_{1}\partial t^{n-1}}|\big)$ for some positive constant $\mathcal{M}$. Recalling Assumption 3, (47), and (49) that renders that $\sigma$ has continuous derivatives of all orders, it means that $\frac{\partial h_{i}}{\partial z_{j}}$, $1\leq j\leq i\leq n$, $i,j\in\mathbb{N}$ are bounded. Applying (48) for $i=2$, recalling (53), we have $z_{2}=\frac{1}{\vartheta}({h_{2}-\frac{\partial h_{1}}{\partial t}-k_{1}h_{1}})$, which is bounded according to the boundedness of $h_{1},h_{2}$, and $\frac{\partial h_{1}}{\partial t}$ which is ensured by the fact that $h_{1}$ is $n$ times differentiable in Assumption 3. Applying (48) for $i=3$, it is obtained that $z_{3}=\frac{1}{\vartheta}({h_{3}-\frac{\partial h_{2}}{\partial t}-k_{2}h_{2}-\frac{\partial h_{2}}{\partial z_{1}}z_{2}})$ is bounded, where $\frac{\partial h_{2}}{\partial t}=\frac{\partial h_{1}^{2}}{\partial z_{1}\partial t}z_{2}+\frac{\partial h_{1}^{2}}{\partial t^{2}}+k_{1}\frac{\partial h_{1}}{\partial t}$ is bounded recalling that $z_{2}$ is bounded as proved above as well as Assumption 3. Similarly, applying (48) for $i=4$, we have that $z_{4}=\frac{1}{\vartheta}({h_{4}-\frac{\partial h_{3}}{\partial t}-k_{3}h_{3}-\frac{\partial h_{3}}{\partial z_{1}}z_{2}-\frac{\partial h_{3}}{\partial z_{2}}z_{3}})$ is bounded, where $\frac{\partial h_{3}}{\partial t}$, which can be expressed by $\frac{\partial h_{1}^{i}}{\partial t^{i}}$ and $z_{i}$, $i=1,2,3$, is bounded according to Assumption 3 and the boundedness of $z_{i}$, $i=1,2,3$ proved above. Recursively applying (48), and repeating the above process until $z_{n}$, we have the inverse transformation of (48) from $H$ to $Z$, and obtain that $|Z(t)|$ is bounded for the boundedness of $|H(t)|$. Recalling (22) and the boundedness of $|v(t)|$ straightforwardly obtained for Sec. 2.2, we have that $|Y(t)|$ is bounded.
According to the theory of Volteral integral equations [38], considering the backstepping transformation (63), (64) as well as the boundedness and continuity of the kernels given in Appendix C, there exists a bounded and continuous the kernel $\mathcal{K}(x,y)$ such that
$\left(\begin{array}[]{c}z(x,t)\\ w(x,t)\\ \end{array}\right)=\left(\begin{array}[]{c}\alpha(x,t)\\ \beta(x,t)\\ \end{array}\right)+\int_{0}^{x}\mathcal{K}(x,y)\left(\begin{array}[]{c}\alpha(y,t)\\ \beta(y,t)\\ \end{array}\right)dy$ (80)
$+\mathcal{K}_{1}(x)Y(t)+\mathcal{K}_{2}(x)v(t)+\left(\begin{array}[]{c}\bar{\varsigma}(x,t)\\ {\varsigma}(x,t)\\ \end{array}\right)$ (83)
$+\int_{0}^{x}\mathcal{K}(x,y)\left(\begin{array}[]{c}\bar{\varsigma}(y,t)\\ {\varsigma}(y,t)\\ \end{array}\right)dy$ (86)
where $\mathcal{K}(x,y)$ can be computed from the kernels ${\phi},\varphi,\Psi,\Phi$, and where $\mathcal{K}_{1}(x)=\left(\gamma(x),\lambda(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\gamma(y),\lambda(y)\right)^{T}dy$, $\mathcal{K}_{2}(x)=\left(\bar{\gamma}(x),\bar{\lambda}(x)\right)^{T}+\int_{0}^{x}\mathcal{K}(x,y)\left(\bar{\gamma}(y),\bar{\lambda}(y)\right)^{T}dy$. Applying the Cauchy-Schwarz inequality to the inverse transformation (86), we have that $z[t],w[t]$ are bounded. Besides, it is obtained from the exponential convergence to zero of $h_{1}$ that $h(e,t)$ is exponentially convergent to zero according to (47) and (49) that shows $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$. Recalling Assumption 3, we know $e(t)$ is exponentially convergent to zero. Property 1 is obtained.
2) According to the choice of the design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ in (73), recalling Lemma 1, we know $h_{i}>0$ at $t=\bar{t}_{0}=\frac{1}{q_{2}}$. Because $\beta(0,t)\equiv 0$ for $t\geq\frac{1}{q_{2}}$ considering the target system (68)–(70), and the system matrix $A_{\rm h}$ defined in (62) represents a form of high-relative-degree CBFs $h_{i}$ as presented in [15], [22], we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Based on this, we prove the safety in cases a)–c) as follows:
a) When the initial safe condition in Definition 1 is satisfied, we know from the fact $h(e(\bar{t}_{0}),\bar{t}_{0})>0$ that $\sigma(t)\equiv 0$ according to (49). Therefore, it is obtained from $h_{1}>0$ for $t\geq\bar{t}_{0}$ and (47) that $h(e(t),t)>0$ for $t\geq\bar{t}_{0}$. Thus, together with fact that $h(e(t),t)\geq 0$ holds on $t\in[t_{0},\bar{t}_{0}]$; ensured by the initial safe condition in Definition 1, we have that $h(e(t),t)\geq 0$ holds on $t\in[t_{0},\infty)$.
b) When the safe initial conditions in Definition 1 is not satisfied, if $h(e(\bar{t}_{0}),\bar{t}_{0})>0$, through the same process with the proof in a), we have $h(e(t),t)\geq 0$ holds on $t\in[\bar{t}_{0},\infty)$;
c) When the safe initial conditions in Definition 1 is not satisfied, if $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, it is followed that $\sigma(t)\equiv 0$ after a finite time $t=\bar{t}_{0}+t_{a}$ according to (49) where $t_{a}>0$ is a free design parameter. We then obtain from the fact $h_{1}>0$ for $t\geq\bar{t}_{0}$ and (47) that $h(e(t),t)\geq 0$ for $t\geq\bar{t}_{0}+t_{a}$. Property 2 is thus obtained. The proof of this theorem is complete.
Remark 3.
If the reference $r(t)\equiv 0$, and the chosen barrier function $h$ satisfying that $\frac{\partial h^{i}(e(t),t)}{\partial t^{i}}$, $i=1,\cdots,n$, converge to zero once $e(t)$ converges to zero, our controller can achieve that full states in the closed-loop system, including the PDE and ODE, are convergent to zero, meanwhile, the output state $C_{1}Y$ stays in the safe region.
4 Unmeasured States and Unknown Disturbances
4.1 State observer and disturbance estimator
In the last section, we have proposed the nominal control design on the basis of a completely known model. In practice, the full states, and also the external disturbances, are always inaccessible. Therefore, we use the measurements $y_{1}(t)=C_{1}Y(t)$, $z(1,t)$, together with the known reference signal $r(t)$, to build an extended observer to estimate the unmeasured states and the unknown external disturbance:
$\dot{\hat{v}}_{r}$ $=S_{r}\hat{v}_{r}(t)+L_{r}(r(t)-\bar{P}_{r}\hat{v}_{r}(t)),$ (87)
$\dot{\hat{v}}_{d}$ $=S_{d}\hat{v}_{d}(t)+L_{d}(z(1,t)-\hat{z}(1,t)),$ (88)
$\dot{\hat{Y}}(t)$ $=A\hat{Y}(t)+B\hat{w}(0,t)+\bar{G}_{1}\hat{v}_{d}(t)+L_{y}(y_{1}(t)-C_{1}\hat{Y}(t))$
$\quad+L_{0}(z(1,t)-\hat{z}(1,t)),$ (89)
${\hat{z}_{t}}(x,t)$ $=-q_{1}{\hat{z}_{x}}(x,t)+{d_{1}}\hat{w}(x,t)+\bar{G}_{2}(x)\hat{v}_{d}(t)$
$\quad+L_{1}(x)(z(1,t)-\hat{z}(1,t)),$ (90)
$\hat{w_{t}}(x,t)$ $=q_{2}\hat{w_{x}}(x,t)+{d_{2}}\hat{z}(x,t)+\bar{G}_{3}(x)\hat{v}_{d}(t)$
$\quad+L_{2}(x)(z(1,t)-\hat{z}(1,t)),$ (91)
$\hat{z}(0,t)$ $=p\hat{w}(0,t)+C\hat{Y}(t)+\bar{G}_{4}\hat{v}_{d}(t),$ (92)
$\hat{w}(1,t)$ $=qz(1,t)+U(t)+\bar{G}_{5}\hat{v}_{d}(t),$ (93)
with the initial states $\hat{z}[t_{0}],\hat{w}[t_{0}],\hat{v}(t_{0})$ chosen in the known bounds given in Assumptions 4, 5, where $\hat{v}_{r},\hat{v}_{d}]$ is the estimate of external signals $v^{T}=[v_{r},v_{d}]$, and $\hat{z}$, $\hat{w}$, $\hat{Y}$ are observer states. The gains $\bar{G}_{i}=G_{i}\bar{P}_{d}$, $i=2,3,4,5$. The observer gain $L_{y}$ is selected such that $A-L_{y}C_{1}$ Hurwitz, considering that $(A,C_{1})$ is observable in Assumption 2. Ohter observer gains $L_{r}$, $L_{y}$, $L_{d}$, $L_{1}(x)$, $L_{2}(x)$ are defined later. Define the estimation error states
[…]
we obtain the observer error system:
$\dot{\tilde{v}}_{r}$ $=(S_{r}-L_{r}\bar{P}_{r})\tilde{v}_{r},$ (95)
$\dot{\tilde{v}}_{d}$ $=S_{d}\tilde{v}_{d}(t)-L_{d}\tilde{z}(1,t),$ (96)
$\dot{\tilde{Y}}(t)$ $=(A-L_{y}C_{1})\tilde{Y}(t)+B\tilde{w}(0,t)+\bar{G}_{1}\tilde{v}_{d}(t)-L_{0}\tilde{z}(1,t)$
${\tilde{z}_{t}}(x,t)$ $=-q_{1}{\tilde{z}_{x}}(x,t)+{d_{1}}\tilde{w}(x,t)$
$\quad+\bar{G}_{2}\tilde{v}_{d}(t)-L_{1}(x)\tilde{z}(1,t),$ (97)
$\tilde{w_{t}}(x,t)$ $=q_{2}\tilde{w_{x}}(x,t)+{d_{2}}\tilde{z}(x,t)$
$\quad-L_{2}(x)\tilde{z}(1,t)+\bar{G}_{3}\tilde{v}_{d}(t),$ (98)
$\tilde{z}(0,t)$ $=p\tilde{w}(0,t)+C\tilde{Y}(t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (99)
$\tilde{w}(1,t)$ $=\bar{G}_{5}\tilde{v}_{d}(t),$ (100)
recalling (1)–(5), (13)–(15). Introduce the transformation:
$\tilde{z}(x,t)=$ $\tilde{\alpha}(x,t)-\int_{x}^{1}K^{11}(x,y)\tilde{\alpha}(y,t)dy$
$-\int_{x}^{1}K^{12}(x,y)\tilde{\beta}(y,t)dy,$ (101)
$\tilde{w}(x,t)=$ $\tilde{\beta}(x,t)-\int_{x}^{1}K^{21}(x,y)\tilde{\alpha}(y,t)dy$
$-\int_{x}^{1}K^{22}(x,y)\tilde{\beta}(y,t)dy$ (102)
${\tilde{Y}(t)=}$ ${\tilde{X}(t)-\int_{0}^{1}K_{0}(x)\tilde{\alpha}(x,t)dx-\int_{0}^{1}K_{1}(x)\tilde{\beta}(x,t)dx}$ (103)
where $K^{11}(x,y)$, $K^{12}(x,y)$, $K^{21}(x,y)$, $K^{22}(x,y)$ on $\{(x,y)|0\leq x\leq y\leq 1\}$ satisfy $q_{1}K^{11}_{x}(x,y)+q_{1}K^{11}_{y}(x,y)=d_{1}K^{21}(x,y)$, $q_{1}K^{12}_{x}(x,y)-q_{2}K^{12}_{y}(x,y)=d_{1}K^{22}(x,y)$, $q_{2}K^{21}_{x}(x,y)-q_{1}K^{21}_{y}(x,y)=-d_{2}K^{11}(x,y)$, $q_{2}K^{22}_{x}(x,y)+q_{2}K^{22}_{y}(x,y)=-d_{2}K^{12}(x,y)$, $K^{21}(x,x)=\frac{-d_{2}}{q_{1}+q_{2}}$, $K^{12}(x,x)=\frac{d_{1}}{q_{1}+q_{2}}$, $K^{11}(0,y)=pK^{21}(0,y)-CK_{0}(y)$, $K^{22}(0,y)=\frac{1}{p}K^{12}(0,y)-\frac{1}{p}CK_{1}(y)$, with $BK^{21}(0,x)-K_{0}^{\prime}(x)q_{1}+(A-L_{y}C_{1})K_{0}(x)=0$, $BK^{22}(0,x)+K_{1}^{\prime}(x)q_{2}+(A-L_{y}C_{1})K_{1}(x)=0$, $B-K_{1}(0)q_{2}=0$, $K_{0}(0)=0$, whose solution can be obtained as the same process in Appendix C (which can be seen clearly by changing the domain from $\{(x,y)|0\leq x\leq y\leq 1\}$ to $\{(x,y)|0\leq y\leq x\leq 1\}$ and comparing it with (B.4)–(B.6) of [30]). Applying the transformations (101)–(103) and choosing the observer gains as
$L_{1}(x)=$ $-p_{1}(x)+\int_{x}^{1}K^{11}(x,y)p_{1}(y)dy$
$+\int_{x}^{1}K^{12}(x,y)p_{2}(y)dy-q_{1}K^{11}(x,1),$ (104)
$L_{2}(x)=$ $-p_{2}(x)+\int_{x}^{1}K^{21}(x,y)p_{1}(y)dy$
$+\int_{x}^{1}K^{22}(x,y)p_{2}(y)dy-q_{1}K^{21}(x,1),$ (105)
where $p_{1}(x),p_{2}(x)$ are to be determined later, then (96)–(100) is converted into
$\dot{\tilde{v}}_{d}$ $=S_{d}\tilde{v}_{d}(t)-L_{d}\tilde{\alpha}(1,t),$ (106)
$\dot{\tilde{X}}(t)$ $=(A-L_{y}C_{1})\tilde{X}(t)+\bar{K}_{0}\tilde{v}_{d}+(\bar{P}_{1}-L_{0})\tilde{\alpha}(1,t)$ (107)
$\tilde{\alpha_{t}}(x,t)$ $=-q_{1}\tilde{\alpha_{x}}(x,t)+\bar{K}_{2}(x)\tilde{v}_{d}(t)+p_{1}(x)\tilde{\alpha}(1,t),$ (108)
$\tilde{\beta_{t}}(x,t)$ $=q_{2}\tilde{\beta_{x}}(x,t)+\bar{K}_{1}(x)\tilde{v}_{d}(t)+p_{2}(x)\tilde{\alpha}(1,t),$ (109)
$\tilde{\beta}(1,t)$ $=\bar{G}_{5}\tilde{v}_{d}(t),$ (110)
$\tilde{\alpha}(0,t)$ $=p\tilde{\beta}(0,t)+C\tilde{X}(t)+\bar{G}_{4}\tilde{v}_{d}(t),$ (111)
where $\bar{P}_{1}=\int_{0}^{1}K_{0}(x)p_{1}(x)dx+\int_{0}^{1}K_{1}(x)p_{2}(x)dx-K_{0}(1)q_{1}$, and where $\bar{K}_{1}(x),\bar{K}_{2}(x)$ satisfy $\bar{K}_{1}(x)-\int_{x}^{1}K^{22}(x,y)\bar{K}_{1}(y)dy$ $\int_{x}^{1}K^{21}(x,y)\bar{K}_{2}(y)dy-q_{2}K^{22}(x,1)\bar{G}_{5}-\bar{G}_{3}(x)=0$, and $\bar{K}_{2}(x)-\int_{x}^{1}K^{11}(x,y)\bar{K}_{2}(y)dy-\int_{x}^{1}K^{12}(x,y)\bar{K}_{1}(y)dy-q_{2}K^{12}(x,1)\bar{G}_{5}-\bar{G}_{2}(x)=0$, and where $\bar{K}_{0}=\int_{0}^{1}K_{0}(x)\bar{K}_{2}(x)dx$ $\bar{G}_{1}+\int_{0}^{1}K_{1}(x)\bar{K}_{1}(x)dx+K_{1}(1)q_{2}\bar{G}_{5}$.
Applying the second transformation:
$\bar{\alpha}(x,t)=$ $\tilde{\alpha}(x,t)-\Lambda(x)\tilde{v}_{d}(t),$ (112)
$\bar{\beta}(x,t)=$ $\tilde{\beta}(x,t)-\Lambda_{1}(x)\tilde{v}_{d}(t),$ (113)
$\tilde{D}(t)=$ $\tilde{X}(t)-\bar{\Lambda}\tilde{v}_{d}(t)$ (114)
where $\Lambda_{1}(x)$, $\Lambda(x)$, $\bar{\Lambda}$ satisfy
$q_{2}\Lambda_{1}^{\prime}(x)-\Lambda_{1}(x)S_{d}+\bar{K}_{1}(x)=0,$ (115)
$\Lambda_{1}(1)=\bar{G}_{5},$ (116)
$q_{1}\Lambda^{\prime}(x)+\Lambda(x)S_{d}-\bar{K}_{2}(x)=0,$ (117)
$\Lambda(0)=p\Lambda_{1}(0)+\bar{G}_{4}+C\bar{\Lambda},$ (118)
$(A-L_{y}C_{1})\bar{\Lambda}+\bar{K}_{0}-\bar{\Lambda}S_{d}=0$ (119)
The solution of (115)–(119) is easy to obtain through the following process. Denote by $\psi_{i}$ the linearly independent eigenvectors of $S_{d}$ associated with the eigenvalues $v_{i}$, $i=1,2,\ldots,n_{d}$. Post-multiplying (115)–(119) by $\psi_{i}$ and introducing
$\iota_{i}=\Lambda\psi_{i},~\iota_{1,i}=\Lambda_{1}\psi_{i},~\bar{\iota}_{i}=\bar{\Lambda}\psi_{i},$ (120)
we obtain $q_{1}\iota^{\prime}_{i}(x)+v_{i}\iota_{i}(x)-\bar{K}_{2}(x)\psi_{i}=0$, $\iota_{i}(0)=p\Lambda_{1}(0)\psi_{i}+\bar{G}_{4}\psi_{i}+C\bar{\Lambda}\psi_{i}$, $q_{2}\iota^{\prime}_{1,i}(x)-v_{i}\iota_{1,i}(x)+\bar{K}_{1}(x)\psi_{i}=0$, $\iota_{1,i}(1)=\bar{G}_{5}\psi_{i}$, $\bar{\iota}_{i}=-((A-L_{y}C_{1})-v_{i}I)^{-1}\bar{K}_{0}\psi_{i}$.
Remark 4.
Since $A-L_{y}C_{1}$ is Hurwitz, all its eigenvalues lie strictly in the open left half-plane. On the other hand, the eigenvalues of $S_{d}$ lie on the imaginary axis, as described in Sec. 2.2. It follows that $\sigma(A-L_{y}C_{1})\cap\sigma(S_{d})=\varnothing.$ Consequently, the matrix $(A-L_{y}C_{1})-v_{i}I$ is invertible.
Therefore, according to (120), the solution of (115)–(119) are $\Lambda=[\iota_{1},\cdots,\iota_{n_{d}}]\bar{V}^{-1}$, $\Lambda_{1}=[\iota_{1,1},\cdots,\iota_{1,n_{d}}]\bar{V}^{-1}$, $\bar{\iota}=[\bar{\iota}_{1},\cdots,\bar{\iota}_{n_{d}}]\bar{V}^{-1}$, where $\bar{V}=[\psi_{1},\cdots,\psi_{n_{d}}]$, and where $\iota_{i}$ and $\iota_{1,i}$ are obtained as
$\iota_{i}(x)=e^{-\frac{v_{i}}{q_{1}}x}\bigg(pe^{\frac{-v_{i}}{q_{2}}}\bar{G}_{5}\psi_{i}+\frac{p}{q_{2}}\int_{0}^{1}e^{\frac{-v_{i}}{q_{2}}\xi}\bar{K}_{1}(\xi)\psi_{i}d\xi$
$+\bar{G}_{4}\psi_{i}-C\big((A-L_{r}C_{1})-v_{i}I\big)^{-1}\bar{K}_{0}\psi_{i}\bigg)$
$+\frac{1}{q_{1}}\int_{0}^{x}e^{-\frac{v_{i}}{q_{1}}(x-\xi)}\bar{K}_{2}(\xi)\psi_{i}d\xi$ (121)
and $\iota_{1,i}(x)=e^{\frac{-v_{i}}{q_{2}}(1-x)}\bar{G}_{5}\psi_{i}-\frac{1}{q_{2}}\int_{x}^{1}e^{\frac{-v_{i}}{q_{2}}(\xi-x)}\bar{K}_{1}(\xi)\psi_{i}d\xi.$
Lemma 2.
The pair $(\Lambda(1),S_{d})$ is observable if and only if $\tilde{N}(v_{i})\bar{P}_{d}\psi_{i}\neq 0$, $i=1,\cdots,n_{d}$, where $\tilde{N}(s)$ is the numerator of the transfer matrix from $d(t)$ to the measurement $z(1,t)$.
Proof.
Applying the same transformation of (101)–(103) (replacing $\tilde{z},\tilde{w},\tilde{Y},\tilde{\alpha},\tilde{\beta},\tilde{X}$ with $z,w,Y,\check{\alpha},\check{\beta},\check{Y}$) into the plant (1)–(5), we obtain $\dot{\check{X}}(t)=(A-L_{r}C_{1})\check{X}(t)-q_{1}K_{0}(1)\check{\alpha}(1,t)+\bar{K}_{0}v_{d}(t),$ $\check{\alpha}_{t}(x,t)=-q_{1}\check{\alpha}_{x}(x,t)-q_{1}K^{11}(x,1)\check{\alpha}(1,t)+\bar{K}_{2}(x)v_{d}(t),$ $\check{\beta}_{t}(x,t)=q_{2}\check{\beta}_{x}(x,t)-q_{1}K^{21}(x,1)\check{\alpha}(1,t)+\bar{K}_{1}(x)v_{d}(t),$ $\check{\beta}(1,t)=q\check{\alpha}(1,t)+\bar{G}_{5}v_{d}(t),$ $\check{\alpha}(0,t)=p\check{\beta}(0,t)+C\check{X}(t)+\bar{G}_{4}v_{d}(t).$ Rewriting it in the frequency domain, considering $z(1,t)=\check{\alpha}(1,t)$, the transfer function from $v_{d}(t)$ to $z(1,t)$ is obtained as $\frac{z(1,s)}{v_{d}(s)}=\frac{N(s)}{D(s)}$, where $D(s)=1+\int_{0}^{1}e^{-(s/q_{1})(1-\xi)}K^{11}(\xi,1)d\xi-pqe^{-s\left(\frac{1}{q_{1}}+\frac{1}{q_{2}}\right)}+e^{-s/q_{1}}p\int_{0}^{1}\frac{q_{1}}{q_{2}}e^{-(s/q_{2})\xi}K^{21}(\xi,1)d\xi+e^{-s/q_{1}}q_{1}C\big(sI-(A-L_{r}C_{1})\big)^{-1}K_{0}(1),$ and where
$N(s)=e^{-s/q_{1}}C\big(sI-(A-L_{r}C_{1})\big)^{-1}\bar{K}_{0}$
$+e^{-s/q_{1}}\Big(p\big(\bar{G}_{5}e^{-s/q_{2}}+\Psi_{1}(s)\big)+\bar{G}_{4}\Big)+\Phi_{2}(s),$ (122)
with $\Phi_{2}(s)=\frac{1}{q_{1}}\int_{0}^{1}e^{-\frac{s}{q_{1}}(1-\xi)}\bar{K}_{2}(\xi)d\xi$ and $\Psi_{1}(s)=\frac{1}{q_{2}}\int_{0}^{1}e^{-\frac{s}{q_{2}}\xi}\bar{K}_{1}(\xi)d\xi.$ Replacing all $\bar{G}_{i}$ involved in $N(s)$ (4.1) with $G_{i}\bar{P}_{d}$, we have $N(s)=\tilde{N}(s)\bar{P}_{d}$, where $\tilde{N}(s)=\frac{z(1,s)}{d(s)}$ is the transfer function from $d(t)$ to $z(1,t)$ considering $d(s)=\bar{P}_{d}v_{d}(s)$. By the PBH (Popov–Belevitch–Hautus) observability test, the pair $(\Lambda(1),S_{d})$ is observable if and only if $\Lambda(1)\psi_{i}\neq 0$, i.e., $\iota_{i}(1)\neq 0$ according to (120). That is, $\iota_{i}(1)=N(v_{i})\psi_{i}=\tilde{N}(v_{i})\bar{P}_{d}\psi_{i}\neq 0$ recalling (4.1).
Remark 5.
Lemma 2 shows the observability condition for the disturbance observer, i.e., all eigenmodes of the disturbance model $\dot{v}_{d}(t)=S_{d}v_{d}(t)$ can be transferred from $d(t)$ to the measurement $z(1,t)$ in steady state, which allows the disturbance observer can reconstruct the state $v_{d}(t)$ from this measurement. Note that $\bar{P}_{d}\psi_{i}\neq 0$ is ensured by Assumption 1 according to PBH observability test.
Defining $p_{1}(x)$, $p_{2}(x)$ in the observer gains (104), (105) are $p_{1}(x)=-\Lambda(x)L_{d}$, $p_{2}(x)=-\Lambda_{1}(x)L_{d}$, then (106)–(111) is converted into
$\dot{\tilde{v}}_{d}$ $=(S_{d}-L_{d}\Lambda(1))\tilde{v}_{d}(t)-L_{d}\bar{\alpha}(1,t),$ (123)
$\dot{\tilde{D}}(t)$ $=(A-L_{y}C_{1})\tilde{D}(t)$ (124)
$\bar{\alpha}_{t}(x,t)$ $=-q_{1}\bar{\alpha}_{x}(x,t),$ (125)
$\bar{\beta}_{t}(x,t)$ $=q_{2}\bar{\beta}_{x}(x,t),$ (126)
$\bar{\beta}(1,t)$ $=0,$ (127)
$\bar{\alpha}(0,t)$ $=p\bar{\beta}(0,t)+C\tilde{D}(t),$ (128)
where $L_{0}$ is chosen as
$L_{0}=\bar{\Lambda}L_{d}+\bar{P}_{1}.$ (129)
The final target observer error system we arrive at is (123)–(128) with (95). The observer gain $L_{d}$ is chosen such that $S_{d}-L_{d}\Lambda(1)$ Hurwitz and $L_{r}$ is selected such that $S_{r}-L_{r}\bar{P}_{r}$ Hurwitz, considering $(S_{d},\Lambda(1))$ and $(S_{r},\bar{P}_{r})$ are observable according to Assumption 1 and Lemma 2.
Before presenting the results, we propose the following lemma and assumptions that will be used in establishing the result about the observer error system. For the observer-based output-feedback control in this section, we require the following assumptions. Assumptions 4, 5 indicate the bounds of the initial values of the unmeasured PDE states and the disturbances are known, but arbitrary, which will be used in estimating the upper bounds of the observer errors required in building the output-feedback safe controller in the next subsection.
Assumption 4.
The bounds of initial states are known, i.e., $\|z(x,0)\|^{2}\leq\overline{z}_{0}$, $\|w(x,0)\|^{2}\leq\overline{w}_{0}$, $|Y(0)|^{2}\leq\overline{Y}_{0}$.
Assumption 5.
The initial values of external signals satisfy $|v(0)|^{2}\in\overline{v}_{0}$, where $\overline{v}_{0}$ is known.
Assumption 6.
The sign of $h(e(\bar{t}_{0}),\bar{t}_{0})$ is known.
For Assumption 6, the exact value of $h(e(\bar{t}_{0}),\bar{t}_{0})$ is not required but its sign that will be used in determining the form of $\sigma$ in (49), which can be judged based on the known bounds of the initial values in Assumptions 4, 5.
Assumption 7.
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}$ is continuous as well as $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[t_{0},\infty)$. Besides, there exists a positive constant $\xi$ such that $|e(t)|<\xi|h(e(t),t)|$. Also, the barrier function $h$ ensures that there exists a positive $\xi_{e}$ such that
$|g(T_{z}Y(t)+T_{v}v(t),z[t],w[t],v(t),t)$
$-g(T_{z}Y(t)+T_{v}\hat{v}(t),\hat{z}[t],\hat{w}[t],\hat{v}(t),t)|$
[…]
where $g$ are determined by $h$ via (72), (54), (47), (48), and (52).
Assumptions 3 and 7 are independent assumptions for the barrier function $h$ in the state-feedback and output-feedback cases, respectively. Due to the uncertainties of the plant state and the external disturbances, Assumption 7 is less general than Assumption 3. Assumption 7 can also be verified easily when planning the barrier function $h$ using (54), (47), (48).
Defining the observer error norm
$\tilde{\Omega}_{e}=(\|\tilde{z}(\cdot,t)\|^{2}+\|\tilde{w}(\cdot,t)\|^{2}+|\tilde{v}|^{2}+|\tilde{Y}|^{2})^{\frac{1}{2}},$ (131)
whose upper bound is estimated in the following lemma, which will be used in building the output-feedback safe controller.
Lemma 3.
For initial data $\hat{w}[0]\in L^{\infty}(0,1)$, $\hat{z}[0]\in L^{\infty}(0,1)$, $\hat{Y}(0)\in\mathbb{R}^{n}$, $\hat{v}_{d}(0)\in\mathbb{R}^{n_{d}}$, $\hat{v}_{r}(0)\in\mathbb{R}^{n_{r}}$, and observability condition in Lemma 2 holds, with the observer (87)–(93) where the observer gains $L_{1}(x)$, $L_{2}(x)$, $L_{0}$ are defined by (104), (105), (129), and $L_{d}$, $L_{r}$, $L_{y}$ satisfies that $S_{d}-L_{d}\Lambda(1)$, $S_{r}-L_{r}\bar{P}_{r}$, and $A-L_{y}C_{1}$ are Hurwitz, respectively, the observer errors in the observer error system (95)–(100) are exponentially convergent to zero in the sense that
$\tilde{\Omega}_{e}\leq\Upsilon_{2}e^{-\sigma_{r}t}:=\rho_{1e}(t),~~~t\in[0,\infty)$ (132)
where $\sigma_{r}>1$ depends on the Hurwitz matrices $S_{d}-L_{d}\Lambda(1)$, $S_{r}-L_{r}\bar{P}_{r}$, $A-L_{y}C_{1}$, and where $\Upsilon_{2}$ depends on the bounds of the kernels in the transformations (101)–(103), (112)–(114), and bounds of the initial state in Assumptions 4 and 5.
Proof.
The estimate (132) can be obtained by the following Lyapunov analysis. Define
$V(t)=$ $\tilde{v}^{T}P_{1}\tilde{v}+{r_{a}}(\tilde{D}^{T}P_{2}\tilde{D}+\frac{b}{2}\int_{0}^{1}\bar{\beta}(x,t)^{2}e^{x}dx)$
$+\frac{r_{c}}{2}\int_{0}^{1}\bar{\alpha}(x,t)^{2}e^{-x}dx$ (133)
where the positive constants $r_{a},r_{b},r_{c}$ are to be determined later. The positive definite matrix $P_{1}={P_{1}}^{T}$ is the solution to the Lyapunov equation $A_{\rm v}^{T}P_{1}+P_{1}A_{\rm v}=-Q_{1}$ for some $Q_{1}={Q_{1}}^{T}>0$, where $A_{\rm v}=[S_{r}-L_{r}\bar{P}_{r},0;0,S_{d}-L_{d}\Lambda(1)]$. The positive definite matrix $P_{2}={P_{2}}^{T}$ is the solution to the Lyapunov equation $(A-L_{y}C_{1})^{T}P_{2}+P_{2}(A-L_{y}C_{1})=-Q_{2}$ for some $Q_{2}={Q_{2}}^{T}>0$. Defining $(\tilde{\Omega}_{T}(t)=\|\bar{\alpha}(\cdot,t)\|^{2}+\|\bar{\beta}(\cdot,t)\|^{2}+|\tilde{v}|^{2}+|\tilde{D}|^{2})^{\frac{1}{2}}$, we have
$\xi_{1}\tilde{\Omega}_{T}(t)^{2}\leq V(t)\leq\xi_{2}\tilde{\Omega}_{T}(t)^{2}$ (134)
for some positive $\xi_{1}$ and $\xi_{2}$. Taking the derivative of (133) along (123)–(128) with (95), applying Young’s and cauthy-schwarz inequalities, we get
$\dot{V}(t)\leq-\frac{1}{2}\lambda_{\rm min}(Q_{1}){{v(t)}^{2}}-\bigg(\frac{{q_{1}}{r_{c}}}{2}-\frac{{2|P_{1}B_{d}|^{2}}}{\lambda_{\rm min}(Q_{1})}\bigg)\bar{\alpha}(1,t)^{2}$
$-\frac{r_{c}}{2}q_{1}\int_{0}^{1}{{}{e^{-x}}{\bar{\alpha}}{{(x,t)}}^{2}}dx-\frac{{r_{a}}q_{2}}{2}\int_{0}^{1}{e^{x}}{\bar{\beta}}{{(x,t)}}^{2}dx$
$-(r_{a}-2r_{c}q_{1}|C|^{2})\lambda_{\rm min}(Q_{2})|\tilde{D}|^{2}$
$-(q_{2}r_{a}-2p^{2}r_{c}q_{1})\bar{\beta}(0,t)^{2}.$ (135)
Choosing $r_{c}>\frac{{4|P_{1}B_{d}|^{2}}}{q_{1}\lambda_{\rm min}(Q_{1})}$, $r_{a}>2r_{c}\max\{2q_{1}|C|^{2},p^{2}\frac{q_{1}}{q_{2}}\}$, where $B_{d}=[0,L_{d}]^{T}$, we arive at $\dot{V}=-\sigma_{v}V(t)$, where $\sigma_{v}=\frac{1}{2\xi_{2}}\min\{\lambda_{\rm min}(Q_{1}),{r_{c}}q_{1}e^{-1},{{r_{a}}q_{2}},{r_{a}}\lambda_{\rm min}(Q_{2})\}$. Therefore, $\tilde{\Omega}_{T}(t)\leq\sqrt{\frac{\xi_{2}}{\xi_{1}}}\tilde{\Omega}_{T}(0)e^{-\sigma_{r}t}$, where $\sigma_{r}=\frac{1}{2}\sigma_{v}$. Estimate (132) is thus obtained by applying the transformations (101)–(103), (112)–(114), and bounding the initial values by Assumptions 4 and 5, i.e., $\tilde{\Omega}_{e}(0)\leq(\overline{z}_{0}+\overline{w}_{0}+\overline{v}_{0}+\overline{Y}_{0})^{\frac{1}{2}}$.
4.2 Output-feedback safe control
Considering the uncertainties, the design parameters are chosen as
$k_{i}$ $>\max_{\ell\in\mathcal{D}_{v}}\{0,\acute{k}_{i}(\ell)\},i=1,2,\cdots,n-1$ (136)
where $\acute{k}_{i}$ are defined in (73), and where $\mathcal{D}_{v}$ denotes the bounds of the disturbances given in Assumption 5. Besides, there also exists conservatism in the choice of $\sigma(t)$, which arises from using known bounds to estimate the lower bound of $h(e(\bar{t}_{0}),\bar{t}_{0})$, denoted as $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})=\min_{\ell_{1},\ell_{2}\in\mathcal{D},\ell_{3}\in\mathcal{D}_{v}}\mathcal{P}(Y(t_{0}),\ell_{1}[t_{0}],\ell_{2}[t_{0}],\ell_{3}(t_{0}))$, where the function $\mathcal{P}$ is defined in (51). The lower bound $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})$ is used in place of $h(e(\bar{t}_{0}),\bar{t}_{0})$ in the coefficient of the function $\sigma(t)$ (49) (the sign of $h(e(\bar{t}_{0}),\bar{t}_{0})$ used in judging the case in (49) is known according to Assumption 6). Design the output-feedback safe controller as
$U_{f}=\hat{U}+{\rm sign}(\vartheta(t_{0}))M_{c}e^{-\sigma_{r}t}$ (137)
where
$\hat{U}=-qz(1,t)+\int_{0}^{1}{\Psi}(1,y)\hat{z}(y,t)dy+\int_{0}^{1}{\Phi}(1,y)\hat{w}(y,t)dy$
$+\lambda(1)\hat{Y}(t)-(\acute{G}_{5}-\bar{\lambda}(1)){\hat{v}}(t)+\hat{\varsigma}(1,t)$ (138)
is obtained by replacing the unmeasurable states $z(x,t)$, $w(x,t)$, $Y(t)$ and and the external disturbances $v(t)$ in the nominal control law (71) with their estimates, and where the sign is defined by ${\rm sign}(u)=\begin{cases}1&u\geq 0\\ -1&u<0\end{cases}$ and $M_{c}$ is given by
$M_{c}=2\max\{\|\Psi(1,y)\|,\|\Phi(1,y)\|,|\acute{G}_{5}+\lambda(1)|,\xi_{e}\}\Upsilon_{2}.$ (139)
The constant $\xi_{e}$ is defined in Assumption 7. The function $\rho_{1e}$ is the upper bound of observer errors, which are given by (132) in Lemma 3. As will be seen clearly later, the function $\rho_{e}(t)$ is indeed an estimated upper bound of the error between the state-feedback and output-feedback controllers. The role of ${\rm sign}(\vartheta(t_{0}))\rho_{e}(t)$ in (137) is to tolerate the estimation errors that arise from the unmeasured states and the external disturbances by shrinking the original safe set to a subset that converges to the original safe set as time approaches infinity.
Theorem 2.
Assume that the closed-loop system consisting of the plant (1)–(5), the observer (87)–(93), and the controller (137) admits a classical solution. For initial data $w[t_{0}]\in L^{\infty}(0,1)$, $z[t_{0}]\in L^{\infty}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, $v(t_{0})\in\mathbb{R}^{n_{v}}$, $\hat{w}[t_{0}]\in L^{\infty}(0,1)$, $\hat{z}[t_{0}]\in L^{\infty}(0,1)$, $\hat{v}(t_{0})\in\mathbb{R}^{n_{v}}$, and for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ satisfying (136), the following properties hold:
1) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
2) Safety is ensured in the sense that:
a) When the safe initial condition in Definition 1 is satisfied, the safety is ensured from $t=t_{0}$, i.e., $h(e(t),t)\geq 0,~\forall t\geq t_{0}$;
b) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})>0$, the safety is ensured from $t=\bar{t}_{0}$, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}$;
c) When the safe initial condition in Definition 1 is not satisfied, and $h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$, the state will return and stay in the safe region no latter than a finite time $\bar{t}_{0}+t_{a}$ where the constant $t_{a}>0$ can be arbitrarily assigned by users, i.e., $h(e(t),t)\geq 0,~\forall t\geq\bar{t}_{0}+t_{a}$.
Proof.
1) Define difference between $U$ and $\hat{U}$ is
$\eta_{e}(t)=$ $U-\hat{U}$
$=$ $-\int_{0}^{1}{\Psi}(1,y)\tilde{z}(y,t)dy-\int_{0}^{1}{\Phi}(1,y)\tilde{w}(y,t)dy$
$-(\acute{G}_{5}+\bar{\lambda}(1))\tilde{v}(t)-\tilde{\varsigma}(1,t),$ (140)
where $\tilde{\varsigma}(1,t)=\varsigma(1,t)-\hat{\varsigma}(1,t)$. Applying the controller (137), we have
$\beta(1,t)={\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))M_{c}e^{-\sigma_{r}t}-\eta_{e}(t).$ (141)
The target system becomes (68)–(69) with (141).
Recalling (139), (140), (72), and (130) in Assumption 7, applying Cauthy-Schwarz inequality and Lemma 3, we obtain
$|\eta_{e}(t)|\leq$ $2\max\{\|\Psi(1,y)\|,\|\Phi(1,y)\|,|\acute{G}_{5}+\lambda(1)|,\xi_{e}\}\tilde{\Omega}_{e}(t)$
$\leq$ $M_{c}e^{-\sigma_{r}t}$ (142)
where $\tilde{\Omega}_{e}(t)$ is defined in (131). Therefore, $\beta(1,t)$ in (141) is exponentially convergent to zero. Through the same process in the proof of the property 1 in Theorem 1, Property 1 is obtained.
2) The sign of $\vartheta$ is the same as its initial value at $t_{0}$ because of the fact that $\vartheta$ is continuous and not zero in Assumption 7 (recalling (52)). It is obtained from (141), (142) and (69), (68) that $\vartheta\beta(0,t)\geq 0$ for $t\in[\bar{t}_{0},\infty)$. The choice of the design parameters $k_{i}$ in (136) is a subset of the ones (73), which thus ensures $h_{i}>0$ at $t=\bar{t}_{0}$ according to Lemma 1. Considering the system matrix $A_{\rm h}$ defined in (62) being a high-relative-degree CBFs $h_{i}$, we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Then, through the same proof as the ones of cases a)–c) of the property 3 in Theorem 1, Property 2 is obtained.
5 Application in UAV Safe Delivery
In the simulation, we apply our control design to a UAV system with cable-suspended payloads for delivery as shown in Fig. 1, aiming to regulate the payload at the lower end of the cable to track a desired reference while avoiding collisions of the payload with surrounding barriers. For simplicity, we consider only one-dimensional motion.
Figure 1: UAV delivery with cable-suspended payloads.
5.1 Model
We adopt the following simplification in modeling: a) We neglect the UAV dynamics, assuming that the control force is applied directly at the top of the cable. However, if the UAV dynamics are taken into account, an additional ODE will be incorporated into the input channel. The corresponding controller can then be designed by combining the results from [30] for ”sandwich” PDE systems. b) Assuming the UAV is flying at a predetermined speed along the $y^{*}$ direction, we only focus on the safe regulation problem in the $x^{*}$ direction. Then, according to [26], the oscillatory dynamics of the cables are modeled by wave PDEs, while the lumped payload at the cable tip is described by a linear ODE. Additionally, wind disturbances $f_{i}$, which act throughout the cable-payload system, are also taken into account. Thus, the mathematical model is formulated as follows:
$M_{0}\ddot{y}_{1}(t)=-d_{0}\dot{y}_{1}(t)+T_{0}u_{x}(0,t)+f_{1}(t),$ (143)
$\rho u_{tt}(x,t)=T_{0}u_{xx}(x,t)-d_{c}{u_{t}}(x,t)+f_{2}(x,t),$ (144)
$u(0,t)=b_{0}(t),~T_{0}u_{x}(L,t)=U(t)+f_{3}(t),$ (145)
for $x\in[0,L]$. Therein, the function $u(x,t)$ represents the distributed transverse displacement along the cable, while $y_{1}(t)$ is the transverse displacement of the payload. The constant $T_{0}=M_{L}g$ denotes the static tension. The terms $f_{i}$ denote airflow disturbances, which will be given in detail later. The values of the physical parameters are listed in Tab. 1. To evaluate the controller under more demanding conditions, we intentionally introduce anti-damping, making the simulation model open-loop unstable. In this setting, instability sources are present in both PDE and ODE subsystems.
Table 1: Physical parameters
Parameters (units) values
Cable length L{L} (m) 1
Cable linear density $\rho$} (kg/m) 0.5
Payload mass $M_{L}$} (kg) 15
Gravitational acceleration $g$ (m/s2) 9.8
[…]
Note: anti-damping is introduced to increase challenges.
Applying Riemann transformations ${z}(x,t)={u_{t}}(x,t)-\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t)$, ${w}(x,t)={u_{t}}(x,t)+\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t)$ and defining a new variable $Y(t)=[y_{1}(t),\dot{y}_{1}(t)]^{T}$, (143)–(145) can be rewritten as
$\dot{Y}(t)=AY(t)+B{w}(0,t)+G_{1}d(t),$ (146)
$z(0,t)=pw(0,t)+CY(t)+G_{4}d(t),$ (147)
${z_{t}}(x,t)=-q_{1}{z_{x}}(x,t)+c{z}(x,t)$
$+c{w}(x,t)+G_{2}(x)d(t),$ (148)
${w_{t}}(x,t)=q_{2}{w_{x}}(x,t)+c{z}(x,t)$
$+c{w}(x,t)+G_{3}(x)d(t),$ (149)
$w(L,t)=qz(L,t)+G_{5}d(t)+U(t),$ (150)
where $q_{1}=q_{2}=\sqrt{\frac{T_{0}}{\rho}},c=\frac{{{-d_{c}}}}{2\rho},p=q=1$, $A=\left[0,1;0,\frac{-d_{0}}{M_{0}}-\frac{\sqrt{T_{0}\rho}}{M_{0}}\right],B=[0,\frac{\sqrt{T_{0}\rho}}{M_{0}}]^{T},C=[0,2].$ After the Riemann transformation in Sec. 5.1, the airflow disturbances $f_{i}$ in the wave PDEs is given by the external disturbances $d(t)=\bar{P}_{d}v_{d}(t)$ (14) with $\bar{P}_{d}=I_{4\times 4}$ and $v_{d}=[\sin(0.25t),\cos(0.25t),\sin(0.5t),\cos(0.5t)]$ in the $2\times 2$ hyperbolic PDEs (146)–(150), where $G_{1}=[g_{1},g_{2}]^{T}$, $g_{1}=[1,0,0,0]$, $g_{2}=[1,1,1,1]$, $G_{4}=[0,1,0,1]$, $G_{5}=[1,0,1,0]$, $G_{2}(x)=[x,0,0,0]$, $G_{2}(x)=[0,x,0,0]$. The desired reference is set as $r(t)=\sin(0.25\pi t)+\cos(0.25\pi t)$, i.e., $\bar{P}_{r}=[1,1]$, and $v_{r}=[\sin(0.25\pi t),\cos(0.25\pi t)]$ in (15). Therefore, the system matrix of exogenous signal model is $S_{d}=[0,0.25,0,0;-0.25,0,0,0;0,0,0,0.5;0,0,$ 0.5,0]-0.5,0] and $S_{r}=[0,0.25\pi;-0.25\pi,0]$.
Remark 6.
Compared to the plant described by (1)–(5), the simulation model represented by (146)–(150) includes additional terms, specifically $c_{1}z(x,t)$ and $c_{2}w(x,t)$ in (148) and (149), respectively. However, these additions do not alter the state-feedback control design presented in this paper. To be exact, applying state-feedback control, the target system defined by (68) and (70), which corresponds to the closed-loop system, remains unchanged except for the updates that (69) are modified to $\beta_{t}=q_{2}\beta_{x}+c\beta$ and $\alpha_{t}=-q_{1}\alpha_{x}+c\alpha$. These modifications clearly do not impact the safety and stability results obtained. The only change in the observer-based output-feedback controller is adding the terms $c\hat{z}(x,t)$ and $c_{2}\hat{w}(x,t)$ in (90) and (91) of the observer, while all other aspects remain unchanged.
The initial values are defined as $w(x,0)=\cos(2\pi x)$, $z(x,0)=\sin(3\pi x)$, $y_{2}(0)=0$, $\hat{w}(x,0)=\cos(2\pi x)+0.5$, $\hat{z}(x,0)=\sin(3\pi x)+0.5$, $\hat{Y}(0)=[\hat{y}_{1}(0),\hat{y}_{2}(0)]^{T}=[0,0]^{T}$, $\hat{v}_{d}(0)=[0.5,1.5,0.5,1.5]^{T}$, $\hat{v}_{r}(0)=[0.5,1.5]^{T}$ where the initial estimation errors are $0.5$. Especially, the values $y_{1}(t_{0})$ will be given in the next subsections. The simulation is conducted using the finite difference method with the space step as $0.05$ and the time space $0.001$.
5.2 Controller and simulation results
We test our safe controller in the following two cases.
5.2.1 Case 1. The safe region is on one side of the desired reference.
The barrier function is set as $h(e(t),t)=e(t)=y_{1}(t)-r(t)$, where $r(t)$ is the desired reference of UAV, and where the region above $r(t)$ is safe. Then we know $\vartheta=\frac{\partial h(e(t),t)}{\partial e}=1$. According to (47)–(49), we have $h_{1}=z_{1}(t)+\sigma(t)$, $h_{2}=z_{2}(t)+\dot{\sigma}(t)+k_{1}z_{1}(t)+k_{1}\sigma(t)$, where $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. According to (54), $f(\underline{z}_{n}(t),t)$ is obtained as, $f(\underline{z}_{n}(t),t)=\frac{1}{b}((k_{1}+k_{2})z_{2}+k_{1}k_{2}z_{1}+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t))$. Then $\varsigma(1,t)$ can be obtained by recalling (72) that uses (B.5). Then the state-feedback controller is obtained recalling (71), where the design parameters are chosen as $k_{1}=1.5$, $k_{2}=4$, and the output-feedback controller $U_{f}$ is obtained applying (137) with choosing the control design parameters $k_{1}=3$, $k_{2}=4$, $M_{c}=15$, $\sigma_{r}=2$, and the observer design parameters $L_{d}=[11,8,11,6]$, $L_{r}=[2,1]$. For safe initialization, we set $y_{1}(t_{0})=8$. For the unsafe initialization case, we set $y_{1}(t_{0})=-1$ that is beyond the safe region, and the design parameters in the output-feedback control law are the same as the output-feedback control in the safe initialization except for $k_{1},k_{2}$ that are chosen as $k_{1}=2.5$, $k_{2}=4.5$ now, and additionally choosing $t_{a}=2$, $\epsilon=2$ in (49) to enforce the out return to the safe region. The simulation results are presented below.
Figure 2: Payload displacement $y_{1}$
Figure 3: Payload velocity $y_{2}$
(a) $w(x,t)$
(b) $z(x,t)$
Figure 4: PDE states with the output-feedback control.
(a) $\tilde{w}(x,t)$
(b) $\tilde{z}(x,t)$
Figure 5: Observer errors
(a) Estimates of $\hat{v}_{d}$ and $\hat{v}_{r}$
(b) Control inputs
Figure 6: Estimates of exogenous signals $v_{d}$, $v_{r}$ given in Sec. 5.1 (red and blue lines are estimates $\hat{v}_{d}$ and $\hat{v}_{r}$, respectively, and dashed black lines are their true values) as well as the control inputs.
We know from the open-loop responses of $y_{1}$ in Fig. 2 that the plant is open-loop unstable, while both state-feedback control and output-feedback control achieve effective safe regulation— tracking the reference and ensuring the position of the payload stays within the safe region. Moreover, when $y_{1}$ starts outside the safe region, the designed control input effectively guides it back to safety within 1.85 seconds, which is less than the designated time $\bar{t}_{0}+t_{a}=2.058$ in the control design. Besides, we know from Fig. 2 that observer state $\hat{y}_{1}$ fast converges to the true value, i.e., the blue dashed line that denotes the evolution of $y_{1}$, whose initial state is in the safe region, under the output-feedback. Another ODE state $y_{2}$, i.e., the payload velocity, is shown in Fig. 3, and the PDE states that reflect the oscillations along the cable are given in Fig. 4 (where we only show the case that the output-feedback control from safe initial states for avoiding repetition), demonstrating the boundedness of these signals. It is also shown in Fig. 3 that the observer state $\hat{y}_{2}$ rapidly converges to its true value $y_{2}$, denoted by the blue dashed line, in the same case (safe initialization). In addition to ODE state estimates, the observer errors for PDE states and the exogenous signals can be seen in Figs 5, 6(a), which show that the observer can successfully track the unmeasured PDE states and the exogenous signals $v_{d}$ and $v_{r}$. The control inputs in the above three situations are given in Fig. 6(b).
5.2.2 Case 2. Barriers are on both sides of the desired reference.
The barrier function is set as
$h(e(t),t)=\delta(t)-|e(t)|=\delta(t)-|y_{1}(t)-r(t)|$ (151)
where the time-varying function $\delta(t)$ is $\delta(t)=15e^{-0.5t}$. It means that the reference trajectory of UAV payload is $r(t)$ and the safe region is within the two barriers: $r(t)+\delta(t)$ and $r(t)-\delta(t)$.
Remark 7.
Because $h$ is not differentiable at $e=0$, the condition in Assumption 7 is not strictly satisfied by (151). However, the controller remains valid by explicitly defining $\frac{\partial h(e,t)}{\partial e}=-1$ when $e=0$, i.e., $\vartheta(e(t),t)=-1$ for $e(t)>0$ and $\vartheta(e(t),t)=1$ for $e(t)\leq 0$. The signal $\vartheta$ is now discontinuous at $e=0$, violating the continuity requirement in Assumption 7. This implies that ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))$ in the output-feedback controller is not guaranteed to be invariant as in the proof of Theorem 2. Nevertheless, with appropriate observer gains $L_{d}$, $L_{r}$ chosen via pole placement, the observer errors can rapidly converge to zero, which means that $\rho_{e}(t)$ decays quickly to zero and the influence of the term ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}$ in the output-feedback controller becomes negligible after a short initial period. Therefore, the proposed control design remains valid without modification, as shown in the upcoming simulation results.
According to (47)–(49), we have $h_{1}=-|z_{1}(t)|+\delta(t)+\sigma(t)$, $h_{2}=\vartheta z_{2}+\dot{\delta}(t)+\dot{\sigma}(t)-k_{1}|z_{1}(t)|+k_{1}\delta(t)+k_{1}\sigma(t)$, with $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. Recalling (54), we obtain $f$ as $f(\underline{z}_{n}(t),t)=\frac{1}{b}\big((k_{1}+k_{2})\vartheta z_{2}-k_{1}k_{2}|z_{1}(t)|+\ddot{\delta}(t)+(k_{1}+k_{2})\dot{\delta}(t)+k_{1}k_{2}\delta(t)+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t)\big)$. Then $\varsigma(1,t)$ can be obtained by recalling (72) that uses (B.5). The output-feedback controller $U_{f}$ is obtained applying (137), (138), where all control design parameters are the same with those in cases 1 except for $k_{1}=30$, $k_{2}=16$, $M_{c}=180$, $\sigma_{r}=1$. For safe initialization, we set $y_{1}(t_{0})=-10$. For unsafe initialization, we set $y_{1}(t_{0})=20$, which falls outside the safe region. The values of control design parameters remain the same in both situations of safe initialization and unsafe initialization. Additionally, we choose $t_{a}=3$ and $\epsilon=2$ in (49) to ensure that the system’s output returns to the safe region in the unsafe initialization situation. The simulation results are presented below.
Figure 7: Payload displacement $y_{1}$
(a) $w(x,t)$
(b) $z(x,t)$
Figure 8: PDE states with the output-feedback control.
It is observed from Fig 7 that when $y_{1}$, i.e., the position of the payload, starts within the safe region, it successfully tracks the reference meanwhile avoiding collisions with the barriers all the time; when the $y_{1}$ starts outside the safe region, it will return to the safe region at 0.34 s, which is much less than $\bar{t}_{0}+t_{a}=2.058$, and also eventually follow the reference. Fig. 8 shows the PDE states are bounded even though anti-damping is added to challenge the task, so is $y_{2}$ whose result is not presented here due to space limit. The results about the observer errors and the estimation errors of the external signals are similar to Figs. 5 and 6, which are not present here for avoiding repetition.
6 Conclusion
This paper has presented a safe output regulation strategy for a class of systems governed by coupled $2\times 2$ hyperbolic PDE-ODE dynamics subject to fully distributed disturbances. By integrating the backstepping method with control barrier function (CBF) techniques, we developed a state-feedback controller that simultaneously enforces safety constraints and ensures exponential convergence of the system output to a desired reference. To address the challenges of unmeasured PDE states and external disturbances, a state observer and disturbance estimator were designed, with explicit bounds on estimation errors incorporated into the control law to ensure robust safety regulation under uncertainty. The resulting output-feedback controller guarantees that the system output remains within a general barrier-defined safe region if the initial condition is safe; otherwise, it is driven back into the safe region within a prescribed time set by the user. All closed-loop signals remain bounded, and the tracking error converges to zero exponentially. The effectiveness of the proposed framework is demonstrated through a UAV delivery scenario with a cable-suspended payload, highlighting its ability to achieve both precise trajectory tracking and reliable collision avoidance under wind disturbances. In future work, safe output regulation control of a class of parabolic PDE-ODE systems will be studied.
Appendix
A The calculation details in the first transformation
Step. 1 Considering the first equation in the equation set (22) with (28), (34), applying (15), we directly have $z_{1}(t)=y_{1}(t)-r(t)=e_{1}(t)$, i.e., (46).
Step. 2 Taking the time derivative of the first equation in the equation set (22) with (28), (34), applying $Y$-ODE (17) and $v$-ODE (12), recalling the second equation in the equation set (22) with (29), (35), (36) for $i=1$, one thus obtains $\dot{z}_{1}(t)=\dot{y}_{1}(t)-P_{r}Sv(t)=a_{11}y_{1}(t)+y_{2}+\acute{g}_{1}v(t)-P_{r}Sv(t)=\varrho_{1,1}y_{1}+y_{2}-(\lambda_{1}+P_{r}S)v(t)=z_{2}(t)$.
Step. 3 Taking the time derivative of the second equation in the equation set (22) with (28), (34), applying $Y$-ODE (17) and $v$-ODE (12), recalling the third equation in the equation set (22) with (30) and (35), (36) for $i=2$, one obtains $\dot{z}_{2}(t)=\dot{y}_{2}(t)+\rho_{11}\dot{y}_{1}(t)-(P_{r}S+\lambda_{1})Sv(t)=a_{2,1}y_{1}+a_{2,2}y_{2}+y_{3}+\acute{g}_{2}v(t)+\rho_{11}a_{11}y_{1}(t)+\rho_{11}y_{2}+\rho_{11}\acute{g}_{1}v(t)-(P_{r}S+\lambda_{1})Sv(t)=\varrho_{2,1}y_{1}+\varrho_{2,2}y_{2}+y_{3}-(\lambda_{2}+P_{r}S^{2})v(t)=z_{3}(t)$.
Step. 4 Similarly, taking the time derivative of the third equation in the equation set (22) with (28), (34), applying $Y$-ODE (17) and $v$-ODE (12), recalling the fourth equation in the equation set (22) with (31)–(33), (35), (36) for $i=3$, one obtains $\dot{z}_{3}(t)=\dot{y}_{3}(t)+\sum_{j=1}^{2}\varrho_{2,j}\dot{y}_{j}-(P_{r}S^{2}+\lambda_{2})Sv(t)=y_{4}+(a_{3,1}+\sum_{j=1}^{2}\varrho_{2,j}a_{j,1})y_{1}+(a_{3,2}+\varrho_{2,1}+\varrho_{2,2}a_{2,2})y_{2}+(a_{3,3}+\varrho_{2,2})y_{3}+(\acute{g}_{3}+\sum_{j=1}^{2}\varrho_{2,j}\acute{g}_{j}-\lambda_{2}S)v(t)-P_{r}S^{3}v(t)=y_{4}+\varrho_{3,1}y_{1}+\varrho_{3,2}y_{2}+\varrho_{3,3}y_{3}-(\lambda_{3}+P_{r}S^{3})v(t)=z_{4}(t)$.
Step. 5 We make an induction hypothesis: for the transformation (22) with (28), (34), we have $\dot{z}_{i}(t)=z_{i+1}(t)$ for $i\geq 3$, under the coefficients $\varrho_{i,\imath},\imath=1,\cdots,i$ and $\lambda_{i}$ defined by (31)–(33), (36), respectively. We then prove that the induction step holds. Taking the time derivative of the $i+1$ th equation in the equation set (22) with (28), (34), applying $Y$-ODE (17) and $v$-ODE (12), we have
$\dot{z}_{i+1}(t)=\dot{y}_{i+1}+\sum_{j=1}^{i}\varrho_{i,j}\dot{y}_{j}-(\lambda_{i}+P_{r}S^{(i)})Sv(t)$
$=y_{i+2}+\sum_{j=1}^{{i+1}}a_{{i+1},j}y_{j}+\sum_{j=2}^{{i+1}}\varrho_{i,j-1}y_{j}+\sum_{j=1}^{i}\sum_{\jmath=1}^{j}\varrho_{i,j}a_{j,\jmath}y_{\jmath}$
$+[\acute{g}_{{i+1}}+\sum_{j=1}^{i}\varrho_{i,j}\acute{g}_{j}-\lambda_{i}S]v(t)-P_{r}S^{(i)}Sv(t)$
$=y_{i+2}+[a_{{i+1},1}+\sum_{j=1}^{i}\varrho_{i,j}a_{j,1}]y_{1}$
$+\sum_{\imath=2}^{i}[a_{{i+1},\imath}+\varrho_{i,\imath-1}+\sum_{j=\imath}^{i}\varrho_{i,j}a_{j,\imath}]y_{\imath}+(a_{{i+1},{i+1}}+\varrho_{i,i})y_{{i+1}}$
$+[\acute{g}_{{i+1}}+\sum_{j=1}^{i}\varrho_{i,j}\acute{g}_{j}-\lambda_{i}S-P_{r}S^{(i+1)}]v(t).$ (A.1)
Recalling $z_{i+2}(t)$ that is defined by the $i+2$ th equation in the equation set (22), for $\dot{z}_{i+1}(t)=z_{i+2}(t)$ to hold, we obtain that the $\varrho_{i+1,\imath},\imath=1,\cdots,i+1$ and $\lambda_{i+1}$ satisfy (31)–(33), (36) for $i+1$. Therefore, the induction step holds. Recalling the Steps 1–4, it follows that $\dot{z}_{i}(t)=z_{i+1}(t)$ hold for $i=1,\cdots,n-1$ via the transformation (22) with (28)–(36).
Step. 6 Taking the time derivative of the $n$ th equation in the equation set (22) with (28), (34), one obtains $\dot{z}_{n}(t)=\dot{y}_{n}+\sum_{j=1}^{n-1}\varrho_{n-1,j}\dot{y}_{j}-(\lambda_{n-1}+P_{r}S^{(n-1)})Sv(t)=\sum_{j=1}^{n}a_{n,j}y_{j}(t)+\acute{g}_{n}v(t)+bw(0,t)+\sum_{j=1}^{n-1}\varrho_{n-1,j}(y_{j+1}+\sum_{\jmath=1}^{j}a_{j,\jmath}y_{\jmath}+\acute{g}_{j}v(t))-(\lambda_{n-1}+P_{r}S^{(n-1)})Sv(t)$. We thus arrive at $\dot{z}_{n}(t)=bw(0,t)+K^{T}Y(t)-[\lambda_{n}+P_{r}S^{n}]v(t)$ with $K^{T}=[\varrho_{n,1},\cdots,\varrho_{n,n}]_{1\times n}$, recalling (31)–(33), (36) for $n$, which is the $n$ th equation the equation set (37) considering (44), (45), and the definition of $B$ and $\bar{G}_{0}$.
B The calculation of the prediction $Z(t+\frac{x}{q_{2}}),Y(t+\frac{x}{q_{2}})$
For (1)–(4), applying (22), and
$\eta(x,t)=w(x,t)-\int_{0}^{x}{\Psi}(x,y)z(y,t)dy$
$-\int_{0}^{x}{\Phi}(x,y)w(y,t)dy-\lambda(x)Y(t)-\bar{\lambda}(x)v(t)$ (B.1)
where the kernels $\Psi$}, $\Phi$}, $\lambda$, $\bar{\lambda}$ are the same as ones that will be given in Appendix C, i.e., (C.16), (C.17) with (C.18)–(C.21), and (C.10), (C.12), (C.13), we then have
$\dot{Z}(t)$ $=A_{z}Z(t)+B\eta(0,t),$ (B.2)
${\eta_{t}}(x,t)$ $=q_{2}{\eta_{x}}(x,t),$ (B.3)
${\eta(1,t)}$ ${=w(1,t)-\int_{0}^{1}{\Psi}(1,y)z(y,t)dy}$
${-\int_{0}^{1}{\Phi}(1,y)w(y,t)dy-\lambda(1)Y(t)-\bar{\lambda}(1)v(t).}$ (B.4)
Therefore, we have
$Z(t+a)=e^{A_{z}a}Z(t)+\int_{0}^{a}e^{A_{z}(a-\tau)}B\eta(q_{2}\tau,t)d\tau,$
for $0\leq a\leq\frac{1}{q_{2}}$, i.e.,
$Z(t+a)=$ $e^{A_{z}a}Z(t)+\frac{1}{q_{2}}\int_{0}^{aq_{2}}e^{A_{z}(a-\frac{\ell}{q_{2}})}B\bigg(w(\ell,t)$
$-\int_{0}^{\ell}{\Psi}(\ell,y)z(y,t)dy-\int_{0}^{\ell}{\Phi}(\ell,y)w(y,t)dy$
$-\lambda(\ell)Y(t)-\bar{\lambda}(\ell)v(t)\bigg)d\ell$
$:=$ $\wp_{1}(Z(t),z[t],w[t],v(t),a),$ (B.5)
for $0\leq a\leq\frac{1}{q_{2}}$. Recalling (12), (22), we have
$Y(t+a)=$ ${T_{z}}^{-1}Z(t+a)-{T_{z}}^{-1}T_{v}v(t+a)$
$=$ ${T_{z}}^{-1}\wp_{1}(T_{z}Y(t)+T_{v}v(t),w[t],z[t],a)$
$-{T_{z}}^{-1}T_{v}e^{Sa}v(t)$
$:=$ $\wp(Y(t),z[t],w[t],v(t),a),~0\leq a\leq\frac{1}{q_{2}},$ (B.6)
recalling (B.3), (B.1). Therefore, $Y(t+\frac{x}{q_{2}})$, $x\in[0,1]$ can be determined by $Y(t)$, $z[t]$, $w[t]$, $v(t)$ as (B.6) with replacing $a$ by $\frac{x}{q_{2}}$.
C Functions $\varphi,\phi,\Psi,\Phi,\gamma,\lambda,\bar{\gamma},\bar{\lambda}$ in (63), (64)
By mapping (37), (2)–(4) and (68)–(69) with using (1), the conditions of the kernels $\varphi,\phi,\Psi,\Phi,\gamma$ and $\lambda$ in the backstepping transformation (63), (64) are obtained as the following equations
${q_{2}}{{\varphi}_{y}}(x,y)-{q_{1}}{{\varphi}_{x}}(x,y)-{d_{1}}\phi(x,y)=0,$ (C.1)
${{q_{1}}{{\phi}_{x}}(x,y){+}{q_{1}}{{\phi}_{y}}(x,y)}{+}{d_{2}}\varphi(x,y)=0,$ (C.2)
${q_{2}}{{\Psi}_{x}}(x,y)-{q_{1}}{{\Psi}_{y}}(x,y)-{d_{2}}\Phi(x,y)=0,$ (C.3)
${{q_{2}}{\Phi_{x}}(x,y)+{q_{2}}{\Phi_{y}}(x,y)-{d_{1}}\Psi(x,y)}=0,$ (C.4)
evolving in the triangular domain $\{(x,y):0\leq y\leq x\leq 1\}$ with the boundary conditions
$\varphi(x,x)=\frac{d_{1}}{{q_{1}}+{q_{2}}},~\Psi(x,x)=\frac{-d_{2}}{{q_{1}}+{q_{2}}},$ (C.5)
$\phi(x,0)=\frac{1}{{q_{1}}p}({q_{2}}\varphi(x,0)-{\gamma}(x)B),$ (C.6)
$\Phi(x,0)=\frac{1}{{q_{2}}}(\lambda(x)B+{q_{1}}p\Psi(x,0)),$ (C.7)
where ${\gamma}(x)$, $\lambda(x)$ satisfy
${q_{1}}{\gamma}^{\prime}(x)+{\gamma}(x)A+{q_{1}}{C}\phi(x,0)=0,$ (C.8)
${q_{2}}\lambda^{\prime}(x)-{\lambda}(x)A-{q_{1}}{C}\Psi(x,0)=0,$ (C.9)
${\lambda(0)=-K^{T}},$ (C.10)
${{\gamma}(0)=-pK^{T}+C},$ (C.11)
with $K^{T}$ defined in (45). The equation set (C.1)–(C.7) is a well-known system of heterodirectional linear coupled hyperbolic PDEs-ODE system, whose well-posedness has been proved in [17]. The functions $\bar{\lambda}$ and $\bar{\gamma}$ satisfy the following two ODEs
$\bar{\lambda}(0)=\frac{1}{b}(\lambda_{n}+P_{r}S^{n}),$ (C.12)
$q_{2}\bar{\lambda}^{\prime}(x)-\bar{\lambda}(x)S-\lambda(x)\acute{G}_{1}+\acute{G}_{3}(x)-\int_{0}^{x}{\Phi}(x,y)\acute{G}_{3}(y)dy$
$-\int_{0}^{x}{\Psi}(x,y)\acute{G}_{2}(y)dy-q_{1}{\Psi}(x,0)\acute{G}_{4}=0,$ (C.13)
$\bar{\gamma}(0)=\frac{p}{b}(\lambda_{n}+P_{r}S^{n})-\acute{G}_{4},$ (C.14)
$-q_{1}\bar{\gamma}^{\prime}(x)-\bar{\gamma}(x)S-\gamma(x)\acute{G}_{1}+\acute{G}_{2}(x)-\int_{0}^{x}{\phi}(x,y)\acute{G}_{2}(y)dy$
$-\int_{0}^{x}{\varphi}(x,y)\acute{G}_{3}(y)dy-q_{1}{\phi}(x,0)\acute{G}_{4}=0.$ (C.15)
The control law requires the solution of $\Psi(x,y)$, $\Phi(x,y)$, whose explicit solution is obtained as follows according to [30] where [23] has been used:
$\Psi(x,y)={F}(x,y)+\int_{y}^{x}L(x,r){F}(r,y)dr,$ (C.16)
$\Phi(x,y)={H}(x,y)-L(x,y)+\int_{y}^{x}L(x,r){H}(r,y)dr,$ (C.17)
where
$F(x,y)=\frac{-1}{p(q_{1}+q_{2})}\bigg[\frac{d_{1}q_{2}}{pq_{1}}I_{0}\left(\frac{2\sqrt{d_{1}d_{2}}}{q_{1}+q_{2}}\sqrt{(x-y)(\frac{q_{1}}{q_{2}}x+y)}\right)$
$+\sqrt{d_{1}d_{2}\frac{x-y}{\frac{q_{1}}{q_{2}}x+y}}I_{1}\left(\frac{2\sqrt{d_{1}d_{2}}}{q_{1}+q_{2}}\sqrt{(x-y)(\frac{q_{1}}{q_{2}}x+y)}\right)$
$+(pd_{2}-\frac{d_{1}q_{2}}{pq_{1}})\Pi\left(\frac{pq_{1}d_{2}}{q_{2}}\frac{x-y}{q_{1}+q_{2}},\frac{d_{1}}{pq_{1}}\frac{q_{1}x+q_{2}y}{q_{1}+q_{2}}\right)\bigg],$ (C.18)
$H(x,y)=\frac{-1}{q_{1}+q_{2}}\bigg[\frac{d_{1}}{p}I_{0}\left(\frac{2\sqrt{d_{1}d_{2}}}{q_{1}+q_{2}}\sqrt{(x-y)(\frac{q_{1}}{q_{2}}x+y)}\right)$
$+\sqrt{d_{1}d_{2}\frac{\frac{q_{1}}{q_{2}}x+y}{x-y}}I_{1}\left(\frac{2\sqrt{d_{1}d_{2}}}{q_{1}+q_{2}}\sqrt{(x-y)(\frac{q_{1}}{q_{2}}x+y)}\right)$
$+(\frac{pd_{2}q_{1}}{q_{2}}-\frac{d_{1}}{p})\Pi\left(\frac{pq_{1}d_{2}}{q_{2}}\frac{x-y}{q_{1}+q_{2}},\frac{d_{1}}{pq_{1}}\frac{q_{1}x+q_{2}y}{q_{1}+q_{2}}\right)\bigg]$ (C.19)
with $I_{j}(j\geq 0)$ denoting the modified Bessel function of the first kind (of order $j$), and $\Pi(s_{1},s_{2})=e^{s_{1}+s_{2}}(1-s_{2}e^{-s_{1}}\int_{0}^{1}e^{-\tau s_{2}}I_{0}(2\sqrt{\tau s_{1}s_{2}})d\tau)$, and where
$L(x,y)=\frac{-1}{q_{2}}\lambda(x-y){B}$ (C.20)
with $\lambda$ given by the ODE
${q_{2}}\lambda^{\prime}(x)-{\lambda}(x)A-{q_{1}}{C}\int_{0}^{x}\frac{-1}{q_{2}}\lambda(z){B}{F}(x-z,0)dz$
$-{q_{1}}{C}{F}(x,0)=0.$ (C.21)